### v1

# Decoding Geometric Properties in Non-Random Data from First Information-Theoretic Principles

Abstract
Based on the principles of information theory, measure theory, and theoretical computer science, we introduce a univariate signal deconvolution method with a wide range of applications to coding theory, particularly in zero-knowledge one-way communication channels, such as in deciphering messages from unknown generating sources about which no prior knowledge is available and to which no return message can be sent. Our multidimensional space reconstruction method from an arbitrary received signal is proven to be agnostic vis-a-vis the encoding-decoding scheme, computation model, programming language, formal theory, the computable (or semi-computable) method of approximation to algorithmic complexity, and any arbitrarily chosen (computable) probability measure of the events. The method derives from the principles of an approach to Artificial General Intelligence capable of building a general-purpose model of models independent of any arbitrarily assumed prior probability distribution. We argue that this optimal and universal method of decoding non-random data has applications to signal processing, causal deconvolution, topological and geometric properties encoding, cryptography, and bio- and technosignature detection.
keywords:
algorithmic information dynamics, causal deconvolution, zero-knowledge communication, technosignatures, signal processing, perturbation analysis, universal distribution, intelligent signal detection, biosignatures
1 Introduction
All of our knowledge about biological systems is currently based on what occurs here on Earth [1, 2, 3, 4, 5, 6, 7, 8, 9]. In order to explore new scenarios in which life and intelligence may appear, in addition to discovering signs or traces of extraterrestrial biological systems, some researchers consider technosignature detection a more promising method of detecting life than other biosignatures because of their longevity [10, 11, 12]. However, some researchers worry that our current theoretical toolkit for data analysis may not be sophisticated enough [13, 14, 15]. Most of the discussion has been centred around the technical justification of hardware choices and the technicalities of detection (e.g. the frequency) and not the nature (i.e., original structural characteristics or underlying mathematical properties) of the signal itself. However, beyond identifying very basic statistical patterns, not much progress has been made on the qualitative semantic aspects of signals and communication, let alone on building a universal framework within which the question can be explored and analysed.
There are myriad data types and formats that encode different types of data on multiple levels, from ASCII characters to video files to health record data. While it is possible to re-encode all data into a single format, the original message may be lost (imagine trying to extract voice-recorded messages of a sound file only by looking at the sound wave plot over time). Because of this diversity even within our own technology, the process of detecting, decoding, and interpreting terrestrial or extraterrestrial signals must involve considering a wide range of possible encoding and decoding schemes.
Without proper conditions that assure the decoding algorithms can only pick up meaningful or sound features—given the prior assumption that a particular dataset actually contains a message—we know the presence of spurious correlations constitutes one of the dangers of naively applying any sort of statistics-based Big Data analysis to extract or find statistically significant patterns in sufficiently large datasets [16, 17]. Ideally, given a random signal such as white noise, no such signal interpretation should yield any meaningful message or interpretation. A random signal should not only look random, but also be interpreted as such, no matter how it is decoded, or how it was encoded. Thus, along with the data compression distortions known to occur from the application of statistics-based or entropy-based methods [18, 19], the possibility of spurious findings in random messages is of major concern in decoding signal streams so as to reconstruct the original message. Unless the receiver agent has prior knowledge about the encoding-decoding scheme beforehand, such an incompressible message would be indistinguishable from an ideal fair-coin-toss sequence of events. Moreover, notice that it is the receiver’s prior knowledge about how to decode the signal according to the emitter’s previously chosen encoding scheme what guarantees the validity of the classical noisy-channel coding theorem [20]. It demonstrates that the emitter agent can introduce redundancies in the original message—thus making the signal data more compressible or non-random—to be correctly interpreted by the receiver so as to avoid decoding noise as meaningful information.
One-way communication channels are those for which the receiver cannot (in principle or in practice) send any signal back to the emitter in order to help or facilitate the decoding process of the first message sent by the emitter. The mathematical phenomena discussed in the above paragraph strongly suggest that for effective communication, especially for one-way communication between any emitter agent and receiver agent, a message should not be displayed in its most compressed form, unless the receiver has knowledge of the compression schema and can uncompress the message back into its original state (disregarding efficiency). Instead, a message should be sent and received in an uncompressed form.
In this case in which the message is encoded in an uncompressed form, or equivalently in which the message is compressible (therefore non-random), one would have a large standard deviation in the range of information values for the different ways of decomposing, parsing or perturbing the signal. Thus, the schema that shortens any redundancies must be read properly by an interpreter to accurately translate the shorthand redundancies back into their longer form. Indeed, it is possible for a signal stream to arrive at a naive receiver in a format that contains sufficient information about how it ought to be decompressed. However, even in this case, the receiver must have prior knowledge about how to decode that signal into its original encoding so as to reveal the redundancies present in the original message that gave it meaning. This prior knowledge means, for example, that one previously knows the chosen encoding-decoding scheme, the programming language, computation model, or the formal theory that explains the relevant physical phenomena which the emitter agent is trying to convey.
In the face of such theoretical challenges, the problem that we tackle in this article is how to encode/decode the message in a compressible signal so as to enable the correct decoding without prior knowledge of the chosen encoding-decoding scheme. We introduce a method based upon the principles of (algorithmic) information theory that is aimed at sweeping over various possible encoding and decoding schemes to test our current limits on signal interpretation as we attempt to reconstruct the original partition (i.e., multidimensional space) in which the original message was given its “meaning” by the emitter agent. This semantic characteristic refers to the (algorithmic) information about the context or real-world correspondents of the emitter that the signal sent by this emitter is trying to convey to the receiver. The real-world correspondents that we particularly investigate are the objects grounded in their respective contexts, that is, embedded into their respective multidimensional spaces as the emitter originally intended.
We demonstrate how messages may be reconstructed by deriving the number of dimensions and the scale of each length of an object from examples ranging from text to images embedded in multiple dimensions, showing a connection between irreducible information content, syntax, geometry, semantics, and topology. This article also shows results that are agnostic vis-a-vis prior knowledge of encoding-decoding schemes, and also demonstrate sufficient conditions in this zero-knowledge scenario for enabling the reconstruction of the original message (i.e., the original object embedded into the original multidimensional space to be conveyed to the receiver) in one-way communication channels.
Zero-knowledge communication occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent, given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of no prior information about the emitter agent only applies to encoding-decoding schemes, and consequentially also to the original multidimensional space and object that are unknown to the receiver before any communication takes place. However, as discussed in the Supplementary Information, this zero-knowledge condition does not mean other assumptions regarding the unknown emitter agent are not being considered by the receiver agent. Also, the reader should not confuse zero-knowledge communication (ZKC) with zero-knowledge proof (ZKP), which is commonly studied in cryptography [21, 22, 23]. Actually, one may consider ZKP and ZKC as akin mathematical problems but as diametrically opposing counterparts regarding the acquisition of knowledge (see Sup. Inf.).
The method is based on the principles of Algorithmic Information Dynamics (AID) [24, 25, 26], and it consists of a perturbation analysis of a received signal. AID is based upon the principles of information theory and the mechanisms of algorithmic probability and the universal distribution, a formal approach to a type of Artificial General/Super Intelligence that requires the massive production of a universal distribution, the mother of all models [27], to build a very large (semi-computable) model of computable approximate models. The underlying idea is that a computable model is a causal explanation of a piece of data. As a result of each perturbation applied to the received signal, a new computable approximate model is built, and then compared against the observation [25, 28].
As explained in Section 4 and formalised in the Sup. Inf., we estimate the algorithmic probability changes for a message under distinct partitions that result from perturbations. This process then builds a large landscape of possible candidate models (along with their respective approximate complexity values) from which one can infer the best model. We show that the partition (i.e., multidimensional space) for which the message displays the lowest complexity indicates the original partition that the received signal stream encodes. For example in the case of bidimensional spaces, our method consists of finding the bidimensional space $\mathcal{S}_{2}$ for which the algorithmic complexity of the object embedded into $\mathcal{S}_{2}$ (i.e., the algorithmic complexity of the message) is minimised in comparison to the landscape of other complexity values calculated from a sufficiently large number of perturbations on the received signal.
The present work applies AID to the problem of communicating messages by performing a perturbation analysis on the received signal. In this manner, as introduced in the Sup. Inf., our present approach also contributes a universal method that is not reducible to a straightforward application of algorithmic information theory (AIT). As the number of perturbed messages strongly dominates the number of possible best message candidates, our theoretical and empirical results show that perturbation analysis enables the downward spikes displayed in the landscapes of complexities to be more likely to indicate the best message candidates, even though the exact values of the algorithmic complexities are semi-computable (i.e., these values are uncomputable and one can only asymptotically approximate them from above). Although there is this intrinsically observer-dependent“deficiency” in the perturbation analysis that forbids an arbitrarily precise approximation to algorithmic complexity, we demonstrate that this subjective aspect can be eventually cancelled out in the perturbation analysis phase. In fact, it is this theoretical limitation given by semi-computability that, counter-intuitively, enables the perturbation analysis to overcome this and the other subjective aspects (such as an arbitrarily chosen programming language, probability measure, or the computational appliances available) involved in the signal processing and interpretation process—that is, in finding an objective piece of information.
2 Results
2.1 On the perturbation analysis of compressed data
Section $4$ of the Sup. Inf. presents empirical results that show our method provides a robust approach to identifying the best way to decode a signal back to its original multidimensional space. Our method consists in sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages by taking the lowest algorithmic complexity configurations under the assumption that the original message is algorithmically not random (i.e., that the original message is compressible). As shown in [29] for graphs, this is because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very low. On the other hand, if the original message is sufficiently compressible (see Section 2.2), and therefore not random, most partitions will lead to configurations of greater algorithmic randomness.
As shown in Figures 11119999 in the Sup. Inf., which apply the method introduced in this article to reconstruct technosignatures, drops (i.e., downward spikes) in the complexity landscapes indicate candidate dimensions for their respective original encoding dimensions, and thus the decoding dimension that would process the signal such that it produces the lowest-complexity message. That is, the original encoding of the message (the one with an interpretable meaning) is the one with the lowest complexity. Our results indicate that sequences that carry meaningful content are those whose complexity quantitatively diverges from the vast majority of the other perturbed messages according to complexity measures. The number of deviations (i.e., perturbations) from the encoded messages that cause the randomness to increase dominates those that do not. In this scenario where a sufficiently large amount of perturbations is taken into account, our results show that an incoming signal with low randomness is more likely to encode a message than one that has more randomness.
In particular, the Block Decomposition Method (BDM) [30] (combining Shannon entropy for long ranges and algorithmic probability for short ranges) was shown to be the complexity measure that was most sensitive and accurate in identifying the original message (image) across the complexity landscapes displayed in Figures $1$, $4$, $6$, $8$, and $9$ in Section $4$ of the Sup. Inf.. Additionally, as shown in Figure $6$, BDM significantly outperforms Compression and Shannon Entropy in the face of additive noise, whose impact on our message reconstruction method is investigated in Figure $5$. As shown in Figures $2$, $3$, and $7$, algorithmic complexity (as estimated via BDM) remains invariant to low-algorithmic-complexity transformations that preserve the algorithmic content of the data (such as swapping all 0s and 1s or swapping all black and white pixels in an image), as well as some non-linear but algorithmic-information-content-preserving transformations [31, 30].
While the above discussed cases study the bidimensional spaces, Fig. LABEL:bone below shows that our method can be scaled up to tridimensional spaces. As we will demonstrate in Section 2.2, we have found sufficient conditions for which the method remains sound for any multidimensional space. The challenges to tackle this generalization is presented in Section $3$ of the Sup. Inf..
2.2 Zero-knowledge one-way communication
The empirical results in this article suggest that AID is capable of harnessing (given the availability of adequate computational resources) some of the algorithmic-informational properties that we demonstrated in ideal and generalised scenarios. As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, most prominently for high algorithmically probable objects.
Our theoretical results show that the method employed to achieve the empirical findings in Section 2.1 is sound with regard to the infinite asymptotic limit when computational resources are unbounded; and it also shows that the method can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Our inference method is general enough to remain robust vis-a-vis any arbitrarily chosen alternative method to calculate the probabilities of the events, any arbitrarily chosen encoding-decoding scheme, programming language, or computation model, without the need of assuming some bias in the possible outcomes and prior beliefs.
Theorem 1 demonstrates that there are conditions for which decoding the message in the original multidimensional space $\mathcal{S}$ chosen from the one in which the message has the lowest algorithmic complexity (i.e., highest algorithmic probability, as assured by the universal distribution) becomes eventually more likely to be correct than decoding the message in a distinct multidimensional space $\mathcal{S}^{\prime}$, with or without noise. The underlying idea is that the larger the algorithmic complexity of the closer-to-randomness message candidate in the wrong multidimensional space with respect to the algorithmic complexity of the message in the correct multidimensional space, the larger the (algorithmic) likelihood [33, 34] of the correct multidimensional space with respect to the likelihood of the wrong closer-to-randomness multidimensional space. Thus, finding the multidimensional space that minimises the algorithmic complexity the most, eventually increases the probability of finding the correct one in comparison to the probability of finding the wrong closer-to-randomness one. Notice that this holds even if one assumes an arbitrarily different probability measure $\mathbf{P}$ as starting point.
We demonstrate that once sufficient conditions for the compressibility of objects, multidimensional spaces, and perturbations (e.g., noise) are met, one-way communication to the receiver becomes possible even if no prior knowledge about the encoding-decoding scheme chosen by the emitter is known. One can first simply look into the landscape of all possible partitions and low-complexity noise the receiver has access to while assuming no bias toward particular objects and partitions in this landscape. Then, despite having no knowledge about the encoding-decoding schemes and thus a lack of biases, one can infer with probability as high as one wishes that the message with the lowest algorithmic complexity is arbitrarily more likely to be correct than any other arbitrary computable method for assigning probabilities to message candidates.
One of the main conditions for this to hold is that the algorithmic complexity of the object $y$ (with or without noise) in the wrong multidimensional space should increase (i.e., move the message toward algorithmic randomness) sufficiently fast in comparison to the algorithmic complexity of the message in a correct multidimensional space. To this end, one can assume a message should contain a large amount of redundancies, and thus should be highly compressible with respect to its linear signal stream length $\left|y\right|$, such as the case in which $\mathbf{K}\left(y\middle|\mathcal{S}\right)=\mathbf{O}\left(\log\log\left(\left|y\right|\right)\right)$ holds, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity. In this manner, most low-complexity perturbations will transform the object and/or partition by sufficiently moving it toward algorithmic randomness so that the conditions are satisfied. This condition resembles the ones for the noisy-channel coding theorem in classical information theory [20], although our method can be understood as being more general than the classical version because: it holds for zero-knowledge communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26]. See also Section $5$ of the Sup. Inf. for further discussion on this and other mathematical phenomena. The proofs of all theorems can be found in Section $3$ of the Sup. Inf..
3 Discussion
We have introduced and demonstrated a general method for reconstructing properties of non-random signals or messages, relating their physical properties to information, entropy, complexity, and semantics. Our method shows how the receiver can decode the (multidimensional) space into which the original message was sent via zero-knowledge one-way communication channels. We have presented signal-amplification techniques that enable the methods to get a better signal-to-noise reading and investigated how the methods perform in the face of noise.
Our method is based on algorithmic information dynamics (AID) which combines tools from classical information theory and algorithmic information theory, in which algorithmic complexity values are approximated and combined with Shannon entropy by using the block decomposition method (BDM), in order to perform a perturbation analysis so that this algorithmic-information-based method can be used to avoid encoding biases in signal detection and processing. Our method is agnostic vis-a-vis endoding-decoding schemes and can be extended and used for bio- and technosignature detection and signal deconvolution. Such results relate information theory to fundamental areas of mathematics, such as geometry and topology, by means of compression and algorithmic probability.
As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM. In addition, although it may seem paradoxical at first glance—because one may initially think that the errors produced by the methods for approximating the algorithmic complexity would be inherited by the perturbation analysis phase itself—our theoretical and empirical results show that the perturbation analysis enables one to overcome the semi-computability limitation in algorithmic information theory, but only when, looking at the whole landscape of complexity values, one estimates with a sufficiently large amount of perturbations effected on the received signal. This is because the algorithmic complexities of the perturbed signals eventually become sufficiently larger than the complexity of the original message. Hence, a powerful enough method for approximating complexity, such as BDM, can eventually harness this divergence in order to enable the perturbation analysis to statistically distinguish the correct one (i.e., more compressible) from the wrong ones (i.e., more random). That is, an approximating method with enough computational resources can produce smaller errors than the difference of actual (although semi-computable) complexity values between the correct candidate and the perturbed ones, even though the exact complexity values remain ultimately unknown to the perturbation analyst (i.e., even though the analyst itself can never be sure it has enough computational resources in the first place). Further discussion of these and other topics can be found in Section $5$ of the Sup. Inf..
4 Methods
4.1 Formalism and basic concepts
For the purposes of the present article, we say an object has “meaning” when the object conveyed in the received signal is actually embedded into (or grounded on) the multidimensional space (i.e., the context) that the emitter agent originally intended. Each (finite and discrete) multidimensional space $\mathcal{S}$—no matter how complex it is—can be univocally determined by the number of dimensions, the encoding of the set of elements for each dimension, and the ordering in which each dimension appears. For example: the unidimensional space $\mathcal{S}_{1}$ that takes values from the natural numbers can be univocally determined by only informing the length $\left|\mathcal{S}_{1}\right|$ in $\mathbf{O}\left(\log\left(\left|\mathcal{S}_{1}\right|\right)\right)$ bits; the bidimensional space $\mathcal{S}_{2}$ by a pair $\left(\left|{\mathcal{S}_{2}}_{x}\right|,\left|{\mathcal{S}_{2}}_{y}\right|\right)$ in $\mathbf{O}\left(\log\left(\left|{\mathcal{S}_{2}}_{x}\right|\right)\right)+\mathbf{O}\left(\log\left(\left|{\mathcal{S}_{2}}_{y}\right|\right)\right)+\mathbf{O}\left(1\right)$ bits, where ${\mathcal{S}_{2}}_{x}$ is the first dimension and ${\mathcal{S}_{2}}_{y}$ is the second dimension, and so on.
In this manner, one is always able to (uniquely) decode the received message into sufficient information to completely determine the multidimensional space configuration. One example of an encoding scheme for an arbitrary multidimensional space $\mathcal{S}$ consists in encoding it in the form of a companion tuple $\bm{\tau}$ as in [35, 36]. Finite and discrete multidimensional spaces with more intricate configurations other than whole number multiplication in the form of $m_{1}\times m_{2}\times\dots\times m_{k}$, where $m,k\in\mathbb{N}$, can, for example, be encoded by node-unaligned companion tuples $\bm{\tau}_{ua}$ as in [36, Definition 2]. Nevertheless, the theoretical framework of our method is agnostic vis-a-vis any arbitrarily chosen computational scheme for encoding multidimensional spaces.
Each partition of a $n$-bit-length linear signal stream corresponds to a distinct configuration of dimension lengths, and therefore to a distinct multidimensional space in which the total additive dimension lengths remains upper bounded by the constant $n$. For example, a picture of 4$\times$4 pixels can be partitioned into other dimension-length configurations leading to different non-squared shapes, like 2$\times$8, 8$\times$2, or 1$\times$16. Once a row (or, alternatively, a column) is fixed in a bidimensional space, the remaining rows and columns are dependent on the initial row fixation. In other words, each bidimensional partition is a single-variable dependency, which is the main rationale underlying the sweep in our algorithms employed to achieve the results in Section 2.1.
More formally, in the case of a (finite and discrete) bidimensional space $\mathcal{S}_{2}$, one may have different combinations of values $\left|{\mathcal{S}_{2}}_{x}\right|$ and $\left|{\mathcal{S}_{2}}_{y}\right|$ for each partition. However, for any partition one has it that both $1\leq\left|{\mathcal{S}_{2}}_{x}\right|\leq n$ and $\lfloor\frac{n}{\left|{\mathcal{S}_{2}}_{x}\right|}\rfloor=\left|{\mathcal{S}_{2}}_{y}\right|$ hold. Thus, because each partition is single-variable dependent on ${\mathcal{S}_{2}}_{x}$, for any $n$-bit-length linear signal stream one can encode each of its $2$D-partitions in $\mathbf{O}\left(\log\left(n\right)\right)$ bits. The same applies to $3$D-partitions (or to any partition from a finite number of dimensions), except for an extra number of bits upper bounded by a partition-independent constant.
[…]
A (algorithmic) perturbation with low algorithmic complexity is one of these transformations where the algorithmic information content of the original object is mostly preserved under such perturbations. More formally, a low-complexity perturbation occurs when $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity, $x$ is the original encoded object, $y$ is the encoded object that results from the perturbation on $x$, and $\left|x\right|$ is the size of the object $x$.
Swapping all 0s for 1s, or swapping all black for white pixels in an image are examples of low-complexity perturbations (in these two cases in particular, one has it that $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(1\right)$). Randomly inserting or deleting a finite number of elements (such as creating or destroying edges in a graph or flipping a finite number of bits in a string) are also examples of low-complexity perturbations [26, 37, 24]. As we saw in the above paragraphs, changing the partition is also a low-complexity perturbation of the multidimensional space into which the object $x$ is embedded. This is because any new partition can always be encoded in $\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$ bits. Notice that a perturbation can change either the object (e.g., by flipping bits) or its multidimensional space (e.g., by reconfiguring the dimension lengths of a given partition to achieve another distinct partition) independently, or can change both at the same time. In any of these three options, except for the identity transformation, the encoded form of the message resulting from the perturbation will be distinct from the encoded form of the original message. Both changing the partition into which the object is embedded and introducing noise into the object itself (e.g., by flipping bits) are particular examples of (algorithmic) perturbations [26, 37, 24]. Notice that these properties of low-complexity perturbations underpin the methodology employed in Section 2.1.
$\mathbf{K}\left(\mathcal{S}\right)$ gives the algorithmic complexity of the arbitrarily chosen encoded form of the multidimensional space $\mathcal{S}$. $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)$ is the algorithmic complexity of the encoded object $y$ once the multidimensional space $\mathcal{S}$ is a priori known. $\mathbf{K}\left(y,\mathcal{S}\right)$ is the algorithmic complexity of the encoding of the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$. Note that there is no loss of generality in our proofs, nor any algorithmic information distortion (except for an independent constant), in employing here the usual notation $\mathbf{K}\left(w,z\right)$, which usually stands for the joint algorithmic complexity of a string $w$ and a string $z$. This is because there is an algorithm that can always encode the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$, if the shortest program that generates the encoding of the pair $\left(y,\mathcal{S}\right)$ is given as input to this algorithm; and there is another algorithm that can always return the encoded pair $\left(y,\mathcal{S}\right)$, if the shortest program that generates the encoded form of the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$ is given as input to this algorithm. For example, as shown in [35, 36], the composite edge set string of a multidimensional network is always computably retrievable from its characteristic string and the encoded form of its companion tuple; and both the characteristic string and the companion tuple of a multidimensional network are computably retrievable from the composite edge set string.
4.2 Toward a general theory for zero-knowledge one-way communication
Let $\mathbf{P}\left[\cdot,\cdot\right]$ be an arbitrarily chosen (computable) probability measure on the space of all possible (computably constructible) finite objects $y$ embedded into a multidimensional space $\mathcal{S}$. For example, $\mathbf{P}\left[y,\mathcal{S}\right]$ may be any probability measure value that is calculated for $y$ embedded into $\mathcal{S}$ by an arbitrarily chosen formal theory $\mathbf{F}$, encoding-decoding scheme, and decoding algorithm.
We know from the algorithmic coding theorem [38, 39, 40, 41] that
$\mathbf{K}\left(x\right)=$ (1)
$-\log\left(\mathbf{P}_{\mathbf{U}}\left[\text{``event }x\text{ occurs''}\right]\right)\pm\mathbf{O}(1)=$
$-\log\left(\mathbf{m}\left(x\right)\right)\pm\mathbf{O}(1)\text{ ,}$
holds, where: $\mathbf{m}\left(\cdot\right)$ is a maximal computably enumerable semimeasure; and $\mathbf{P}_{\mathbf{U}}\left[\cdot\right]$ denotes the universal a priori probability of an arbitrary event. $\mathbf{P}_{\mathbf{U}}$ can be understood as the probability of randomly generating (by an i.i.d. stochastic process) a prefix-free (or self-delimiting) program that outputs $x$. A computably enumerable semimeasure $\mathbf{m}\left(\cdot\right)$ is said to be maximal if, for any other computably enumerable semimeasure $\mu\left(\cdot\right)$ with domain defined for possible encoded objects, where $\sum\limits_{x\in\left\{0,1\right\}^{*}}\mu\left(x\right)\leq 1$, there is a constant $C>0$ (which does not depend on $x$) such that, for every encoded object $x$, $\mathbf{m}\left(x\right)\geq C\,\mu\left(x\right)\text{ .}$ Also note that the algorithmic coding theorem applies analogously to the conditional algorithmic complexity $\mathbf{K}(z\,|w)$.
We can explore the power and universality of the algorithmic coding theorem in order to investigate a lower bound for the universal (probability) distribution in comparison to an arbitrary probability measure $\mathbf{P}$. A chosen probability measure $\mathbf{P}$ may, or may not, assign a higher probability value to a correct multidimensional space. Thus, it is important to evaluate how much a universal approach such as the universal a priori probability can be wrong (i.e., diverges from $\mathbf{P}$) in the asymptotic limit. Indeed, the algorithmic coding theorem enables one to guarantee that the multiplicative error of
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]=\mathbf{P}_{\mathbf{U}}\left[y\middle|\mathcal{S}\right]$
with respect to
$\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]$
is bounded by the algorithmic probability of the inverse algorithmic perturbation $\mathcal{P}^{-1}$ itself. This means that the easier, or less complex, it is to retrieve $y$ embedded into $\mathcal{S}$ from $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$ the smaller the multiplicative error of employing $\mathbf{P}_{\mathbf{U}}$ instead of the arbitrarily chosen $\mathbf{P}$.
In the general case (where the multidimensional spaces may be much more complex than the objects themselves), even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended [35, 36]. (See also Section $2$ in the Sup. Inf..) This poses a challenge to straightforward applications of the algorithmic information theory (AIT). In order to tackle such a challenge, we develop a method for multidimensional reconstruction of received low-dimensional signals in which one does not have access to knowledge about the encoding-decoding scheme or the original multidimensional space. Theorem 1 demonstrates the existence of sufficient conditions (or assumptions) that allow one to achieve an encoding/decoding-agnostic method for reconstructing the original multidimensional spaces. It gives rise to an agnostic method that does not need to assume that one a priori knows that the emitter agent’s multidimensional space resembles ours, is as complex as ours, or is isomorphic to the structures we currently deem fundamental in mathematics.
Theorem 1 (Theorem $3.10$ in the Sup. Inf.).
Let $y$ be an arbitrary object embedded into an arbitrary multidimensional space $\mathcal{S}$ such that the following conditions
  • $\mathbf{K}\left(\mathcal{S}\right)=\mathbf{o}\big{(}\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\big{)}\text{ ,}$ (2)
  • $\mathbf{K}\left(\mathcal{P}\right)=\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}\text{ ,}$ (3)
  • and
  • $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\leq\mathbf{o}\big{(}\log\left(\left|y\right|\right)\big{)}$ (4)
hold, where $\mathcal{P}$ is any algorithmic perturbation that transforms $y$ embedded into $\mathcal{S}$ into another $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$. Let $\mathbf{P}^{\prime}\left[\cdot,\cdot\right]$ be the uniform probability measure over the space of all possible outcomes (in the form $\left(y^{\prime},\,\mathcal{S}^{\prime}\right)$) of algorithmic perturbations on $\left(y,\,\mathcal{S}\right)$ that satisfy Equation 3. Let $\mathbf{P}\left[\cdot,\cdot\right]$ be any arbitrarily chosen (computable) probability measure on the space of all possible (computably constructible) objects embedded into a multidimensional space. Then, the following strong asymptotic dominance
$\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]=\mathbf{o}\big{(}\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]\big{)}$ ” ] ) (5)
holds with probability (given by the uniform probability measure $\mathbf{P}^{\prime}$) arbitrarily close to $1$ as $\left|y\right|\to\infty$ for $\left(y^{\prime},\mathcal{S}^{\prime}\right)$ distinct from $\left(y,\mathcal{S}\right)$ satisfying Equation 3.
By applying AID to the problem of communicating messages so that one can perform a perturbation analysis on the received signal, Theorem 1 enables one to exploit the semi-computability of algorithmic complexity while taking advantage of the divergence brought about by the landscape of the approximate complexity values, landscape which results from the perturbations and from which we can find the best candidates to choose. Individually, the approximation to the complexity value of each object (and multidimensional space) is prone to the unknown error margins that follow from its semi-computability. However, because each pair object-space results from a perturbation, which is a particular case of algorithm, and each algorithm theoretically has an exact complexity value associated to it, one can retrieve a likely correct answer from a sufficiently large landscape of these perturbations.
As further discussed in Section $5$ of the Sup. Inf., we believe the results in this article not only contribute with a general method for interpreting messages in zero-knowledge one-way communication channels, but also establish fundamental connections (and posit open problems) with zero-knowledge proofs and the algorithmic counterparts of the maximum likelihood method and the noisy-channel coding theorem.
[…]
Acknowledgements
Felipe S. Abrahão acknowledges support from the São Paulo Research Foundation (FAPESP), grants $2021$/$14501$-$8$ and $2023$/$05593$-$1$.