### v2

# Decoding Geometric Properties in Non-Random Data from First Information-Theoretic Principles

Abstract
Based on the principles of information theory, measure theory, and theoretical computer science, we introduce a univariate signal deconvolution method with a wide range of applications to coding theory, particularly in zero-knowledge one-way communication channels, such as in deciphering messages from unknown generating sources about which no prior knowledge is available and to which no return message can be sent. Our multidimensional space reconstruction method from an arbitrary received signal is proven to be agnostic vis-a-vis the encoding-decoding scheme, computation model, programming language, formal theory, the computable (or semi-computable) method of approximation to algorithmic complexity, and any arbitrarily chosen (computable) probability measure of the events. The method derives from the principles of an approach to Artificial General Intelligence capable of building a general-purpose model of models independent of any arbitrarily assumed prior probability distribution. We argue that this optimal and universal method of decoding non-random data has applications to signal processing, causal deconvolution, topological and geometric properties encoding, cryptography, and bio- and technosignature detection.
keywords:
algorithmic information dynamics, causal deconvolution, zero-knowledge communication, technosignatures, signal processing, perturbation analysis, universal distribution, intelligent signal detection, biosignatures
1 Introduction
All of our knowledge about biological systems is currently based on what occurs here on Earth [1, 2, 3, 4, 5, 6, 7, 8, 9]. In order to explore new scenarios in which life and intelligence may appear, in addition to discovering signs or traces of extraterrestrial biological systems, some researchers consider technosignature detection a more promising method of detecting life than other biosignatures because of their longevity [10, 11, 12]. However, some researchers worry that our current theoretical toolkit for data analysis may not be sophisticated enough [13, 14, 15]. Most of the discussion has been centred around the technical justification of hardware choices and the technicalities of detection (e.g. the frequency) and not the nature (i.e., original structural characteristics or underlying mathematical properties) of the signal itself. However, beyond identifying very basic statistical patterns, not much progress has been made on the qualitative semantic aspects of signals and communication, let alone on building a universal framework within which the question can be explored and analysed.
There are myriad data types and formats that encode different types of data on multiple levels, from ASCII characters to video files to health record data. While it is possible to re-encode all data into a single format, the original message may be lost (imagine trying to extract voice-recorded messages of a sound file only by looking at the sound wave plot over time). Because of this diversity even within our own technology, the process of detecting, decoding, and interpreting terrestrial or extraterrestrial signals must involve considering a wide range of possible encoding and decoding schemes.
Without proper conditions that assure the decoding algorithms can only pick up meaningful or sound features—given the prior assumption that a particular dataset actually contains a message—we know the presence of spurious correlations constitutes one of the dangers of naively applying any sort of statistics-based Big Data analysis to extract or find statistically significant patterns in sufficiently large datasets [16, 17]. Ideally, given a random signal such as white noise, no such signal interpretation should yield any meaningful message or interpretation. A random signal should not only look random, but also be interpreted as such, no matter how it is decoded, or how it was encoded. Thus, along with the data compression distortions known to occur from the application of statistics-based or entropy-based methods [18, 19], the possibility of spurious findings in random messages is of major concern in decoding signal streams so as to reconstruct the original message. Unless the receiver agent has prior knowledge about the encoding-decoding scheme beforehand, such an incompressible message would be indistinguishable from an ideal fair-coin-toss sequence of events. Moreover, notice that it is the receiver’s prior knowledge about how to decode the signal according to the emitter’s previously chosen encoding scheme what guarantees the validity of the classical noisy-channel coding theorem [20]. It demonstrates that the emitter agent can introduce redundancies in the original message—thus making the signal data more compressible or non-random—to be correctly interpreted by the receiver so as to avoid decoding noise as meaningful information.
One-way communication channels are those for which the receiver cannot (in principle or in practice) send any signal back to the emitter in order to help or facilitate the decoding process of the first message sent by the emitter. The mathematical phenomena discussed in the above paragraph strongly suggest that for effective communication, especially for one-way communication between any emitter agent and receiver agent, a message should not be displayed in its most compressed form, unless the receiver has knowledge of the compression schema and can uncompress the message back into its original state (disregarding efficiency). Instead, a message should be sent and received in an uncompressed form.
In this case in which the message is encoded in an uncompressed form, or equivalently in which the message is compressible (therefore non-random), one would have a large standard deviation in the range of information values for the different ways of decomposing, parsing or perturbing the signal. Thus, the schema that shortens any redundancies must be read properly by an interpreter to accurately translate the shorthand redundancies back into their longer form. Indeed, it is possible for a signal stream to arrive at a naive receiver in a format that contains sufficient information about how it ought to be decompressed. However, even in this case, the receiver must have prior knowledge about how to decode that signal into its original encoding so as to reveal the redundancies present in the original message that gave it meaning. This prior knowledge means, for example, that one previously knows the chosen encoding-decoding scheme, the programming language, computation model, or the formal theory that explains the relevant physical phenomena which the emitter agent is trying to convey.
In the face of such theoretical challenges, the problem that we tackle in this article is how to encode/decode the message in a compressible signal so as to enable the correct decoding without prior knowledge of the chosen encoding-decoding scheme. We introduce a method based upon the principles of (algorithmic) information theory that is aimed at sweeping over various possible encoding and decoding schemes to test our current limits on signal interpretation as we attempt to reconstruct the original partition (i.e., multidimensional space) in which the original message was given its “meaning” by the emitter agent. This semantic characteristic refers to the (algorithmic) information about the context or real-world correspondents of the emitter that the signal sent by this emitter is trying to convey to the receiver. The real-world correspondents that we particularly investigate are the objects grounded in their respective contexts, that is, embedded into their respective multidimensional spaces as the emitter originally intended.
We demonstrate how messages may be reconstructed by deriving the number of dimensions and the scale of each length of an object from examples ranging from text to images embedded in multiple dimensions, showing a connection between irreducible information content, syntax, geometry, semantics, and topology. This article also shows results that are agnostic vis-a-vis prior knowledge of encoding-decoding schemes, and also demonstrate sufficient conditions in this zero-knowledge scenario for enabling the reconstruction of the original message (i.e., the original object embedded into the original multidimensional space to be conveyed to the receiver) in one-way communication channels.
Zero-knowledge communication occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent, given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of no prior information about the emitter agent only applies to encoding-decoding schemes, and consequentially also to the original multidimensional space and object that are unknown to the receiver before any communication takes place. However, as discussed in the Supplementary Information, this zero-knowledge condition does not mean other assumptions regarding the unknown emitter agent are not being considered by the receiver agent. Also, the reader should not confuse zero-knowledge communication (ZKC) with zero-knowledge proof (ZKP), which is commonly studied in cryptography [21, 22, 23]. Actually, one may consider ZKP and ZKC as akin mathematical problems but as diametrically opposing counterparts regarding the acquisition of knowledge (see Sup. Inf.).
The method is based on the principles of Algorithmic Information Dynamics (AID) [24, 25, 26], and it consists of a perturbation analysis of a received signal. AID is based upon the principles of information theory and the mechanisms of algorithmic probability and the universal distribution, a formal approach to a type of Artificial General/Super Intelligence that requires the massive production of a universal distribution, the mother of all models [27], to build a very large (semi-computable) model of computable approximate models. The underlying idea is that a computable model is a causal explanation of a piece of data. As a result of each perturbation applied to the received signal, a new computable approximate model is built, and then compared against the observation [25, 28].
As explained in Section 4 and formalised in the Sup. Inf., we estimate the algorithmic probability changes for a message under distinct partitions that result from perturbations. This process then builds a large landscape of possible candidate models (along with their respective approximate complexity values) from which one can infer the best model. We show that the partition (i.e., multidimensional space) for which the message displays the lowest complexity indicates the original partition that the received signal stream encodes. For example in the case of bidimensional spaces, our method consists of finding the bidimensional space $\mathcal{S}_{2}$ for which the algorithmic complexity of the object embedded into $\mathcal{S}_{2}$ (i.e., the algorithmic complexity of the message) is minimised in comparison to the landscape of other complexity values calculated from a sufficiently large number of perturbations on the received signal.
The present work applies AID to the problem of communicating messages by performing a perturbation analysis on the received signal. In this manner, as introduced in the Sup. Inf., our present approach also contributes a universal method that is not reducible to a straightforward application of algorithmic information theory (AIT). As the number of perturbed messages strongly dominates the number of possible best message candidates, our theoretical and empirical results show that perturbation analysis enables the downward spikes displayed in the landscapes of complexities to be more likely to indicate the best message candidates, even though the exact values of the algorithmic complexities are semi-computable (i.e., these values are uncomputable and one can only asymptotically approximate them from above). Although there is this intrinsically observer-dependent“deficiency” in the perturbation analysis that forbids an arbitrarily precise approximation to algorithmic complexity, we demonstrate that this subjective aspect can be eventually cancelled out in the perturbation analysis phase. In fact, it is this theoretical limitation given by semi-computability that, counter-intuitively, enables the perturbation analysis to overcome this and the other subjective aspects (such as an arbitrarily chosen programming language, probability measure, or the computational appliances available) involved in the signal processing and interpretation process—that is, in finding an objective piece of information.
2 Results
2.1 On the perturbation analysis of compressed data
Section A.4 presents empirical results that show our method provides a robust approach to identifying the best way to decode a signal back to its original multidimensional space. Our method consists in sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages by taking the lowest algorithmic complexity configurations under the assumption that the original message is algorithmically not random (i.e., that the original message is compressible). As shown in [29] for graphs, this is because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very low. On the other hand, if the original message is sufficiently compressible (see Section 2.2), and therefore not random, most partitions will lead to configurations of greater algorithmic randomness.
As shown in Figures 11119999 in the Sup. Inf., which apply the method introduced in this article to reconstruct technosignatures, drops (i.e., downward spikes) in the complexity landscapes indicate candidate dimensions for their respective original encoding dimensions, and thus the decoding dimension that would process the signal such that it produces the lowest-complexity message. That is, the original encoding of the message (the one with an interpretable meaning) is the one with the lowest complexity. Our results indicate that sequences that carry meaningful content are those whose complexity quantitatively diverges from the vast majority of the other perturbed messages according to complexity measures. The number of deviations (i.e., perturbations) from the encoded messages that cause the randomness to increase dominates those that do not. In this scenario where a sufficiently large amount of perturbations is taken into account, our results show that an incoming signal with low randomness is more likely to encode a message than one that has more randomness.
In particular, the Block Decomposition Method (BDM) [30] (combining Shannon entropy for long ranges and algorithmic probability for short ranges) was shown to be the complexity measure that was most sensitive and accurate in identifying the original message (image) across the complexity landscapes displayed in Figures $1$, $4$, $6$, $8$, and $9$ in Section A.4. Additionally, as shown in Figure $6$, BDM significantly outperforms Compression and Shannon Entropy in the face of additive noise, whose impact on our message reconstruction method is investigated in Figure $5$. As shown in Figures $2$, $3$, and $7$, algorithmic complexity (as estimated via BDM) remains invariant to low-algorithmic-complexity transformations that preserve the algorithmic content of the data (such as swapping all 0s and 1s or swapping all black and white pixels in an image), as well as some non-linear but algorithmic-information-content-preserving transformations [31, 30].
While the above discussed cases study the bidimensional spaces, Fig. LABEL:bone below shows that our method can be scaled up to tridimensional spaces. As we will demonstrate in Section 2.2, we have found sufficient conditions for which the method remains sound for any multidimensional space. The challenges to tackle this generalization is presented in Section A.2.
2.2 Zero-knowledge one-way communication
The empirical results in this article suggest that AID is capable of harnessing (given the availability of adequate computational resources) some of the algorithmic-informational properties that we demonstrated in ideal and generalised scenarios. As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, most prominently for high algorithmically probable objects.
Our theoretical results show that the method employed to achieve the empirical findings in Section 2.1 is sound with regard to the infinite asymptotic limit when computational resources are unbounded; and it also shows that the method can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Our inference method is general enough to remain robust vis-a-vis any arbitrarily chosen alternative method to calculate the probabilities of the events, any arbitrarily chosen encoding-decoding scheme, programming language, or computation model, without the need of assuming some bias in the possible outcomes and prior beliefs.
Theorem 1 demonstrates that there are conditions for which decoding the message in the original multidimensional space $\mathcal{S}$ chosen from the one in which the message has the lowest algorithmic complexity (i.e., highest algorithmic probability, as assured by the universal distribution) becomes eventually more likely to be correct than decoding the message in a distinct multidimensional space $\mathcal{S}^{\prime}$, with or without noise. The underlying idea is that the larger the algorithmic complexity of the closer-to-randomness message candidate in the wrong multidimensional space with respect to the algorithmic complexity of the message in the correct multidimensional space, the larger the (algorithmic) likelihood [33, 34] of the correct multidimensional space with respect to the likelihood of the wrong closer-to-randomness multidimensional space. Thus, finding the multidimensional space that minimises the algorithmic complexity the most, eventually increases the probability of finding the correct one in comparison to the probability of finding the wrong closer-to-randomness one. Notice that this holds even if one assumes an arbitrarily different probability measure $\mathbf{P}$ as starting point.
We demonstrate that once sufficient conditions for the compressibility of objects, multidimensional spaces, and perturbations (e.g., noise) are met, one-way communication to the receiver becomes possible even if no prior knowledge about the encoding-decoding scheme chosen by the emitter is known. One can first simply look into the landscape of all possible partitions and low-complexity noise the receiver has access to while assuming no bias toward particular objects and partitions in this landscape. Then, despite having no knowledge about the encoding-decoding schemes and thus a lack of biases, one can infer with probability as high as one wishes that the message with the lowest algorithmic complexity is arbitrarily more likely to be correct than any other arbitrary computable method for assigning probabilities to message candidates.
One of the main conditions for this to hold is that the algorithmic complexity of the object $y$ (with or without noise) in the wrong multidimensional space should increase (i.e., move the message toward algorithmic randomness) sufficiently fast in comparison to the algorithmic complexity of the message in a correct multidimensional space. To this end, one can assume a message should contain a large amount of redundancies, and thus should be highly compressible with respect to its linear signal stream length $\left|y\right|$, such as the case in which $\mathbf{K}\left(y\middle|\mathcal{S}\right)=\mathbf{O}\left(\log\log\left(\left|y\right|\right)\right)$ holds, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity. In this manner, most low-complexity perturbations will transform the object and/or partition by sufficiently moving it toward algorithmic randomness so that the conditions are satisfied. This condition resembles the ones for the noisy-channel coding theorem in classical information theory [20], although our method can be understood as being more general than the classical version because: it holds for zero-knowledge communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26]. See also Section A.5 for further discussion on this and other mathematical phenomena. The proofs of all theorems can be found in Section 2.2.
3 Discussion
We have introduced and demonstrated a general method for reconstructing properties of non-random signals or messages, relating their physical properties to information, entropy, complexity, and semantics. Our method shows how the receiver can decode the (multidimensional) space into which the original message was sent via zero-knowledge one-way communication channels. We have presented signal-amplification techniques that enable the methods to get a better signal-to-noise reading and investigated how the methods perform in the face of noise.
Our method is based on algorithmic information dynamics (AID) which combines tools from classical information theory and algorithmic information theory, in which algorithmic complexity values are approximated and combined with Shannon entropy by using the block decomposition method (BDM), in order to perform a perturbation analysis so that this algorithmic-information-based method can be used to avoid encoding biases in signal detection and processing. Our method is agnostic vis-a-vis endoding-decoding schemes and can be extended and used for bio- and technosignature detection and signal deconvolution. Such results relate information theory to fundamental areas of mathematics, such as geometry and topology, by means of compression and algorithmic probability.
As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM. In addition, although it may seem paradoxical at first glance—because one may initially think that the errors produced by the methods for approximating the algorithmic complexity would be inherited by the perturbation analysis phase itself—our theoretical and empirical results show that the perturbation analysis enables one to overcome the semi-computability limitation in algorithmic information theory, but only when, looking at the whole landscape of complexity values, one estimates with a sufficiently large amount of perturbations effected on the received signal. This is because the algorithmic complexities of the perturbed signals eventually become sufficiently larger than the complexity of the original message. Hence, a powerful enough method for approximating complexity, such as BDM, can eventually harness this divergence in order to enable the perturbation analysis to statistically distinguish the correct one (i.e., more compressible) from the wrong ones (i.e., more random). That is, an approximating method with enough computational resources can produce smaller errors than the difference of actual (although semi-computable) complexity values between the correct candidate and the perturbed ones, even though the exact complexity values remain ultimately unknown to the perturbation analyst (i.e., even though the analyst itself can never be sure it has enough computational resources in the first place). Further discussion of these and other topics can be found in Section A.5.
4 Methods
4.1 Formalism and basic concepts
For the purposes of the present article, we say an object has “meaning” when the object conveyed in the received signal is actually embedded into (or grounded on) the multidimensional space (i.e., the context) that the emitter agent originally intended. Each (finite and discrete) multidimensional space $\mathcal{S}$—no matter how complex it is—can be univocally determined by the number of dimensions, the encoding of the set of elements for each dimension, and the ordering in which each dimension appears. For example: the unidimensional space $\mathcal{S}_{1}$ that takes values from the natural numbers can be univocally determined by only informing the length $\left|\mathcal{S}_{1}\right|$ in $\mathbf{O}\left(\log\left(\left|\mathcal{S}_{1}\right|\right)\right)$ bits; the bidimensional space $\mathcal{S}_{2}$ by a pair $\left(\left|{\mathcal{S}_{2}}_{x}\right|,\left|{\mathcal{S}_{2}}_{y}\right|\right)$ in $\mathbf{O}\left(\log\left(\left|{\mathcal{S}_{2}}_{x}\right|\right)\right)+\mathbf{O}\left(\log\left(\left|{\mathcal{S}_{2}}_{y}\right|\right)\right)+\mathbf{O}\left(1\right)$ bits, where ${\mathcal{S}_{2}}_{x}$ is the first dimension and ${\mathcal{S}_{2}}_{y}$ is the second dimension, and so on.
In this manner, one is always able to (uniquely) decode the received message into sufficient information to completely determine the multidimensional space configuration. One example of an encoding scheme for an arbitrary multidimensional space $\mathcal{S}$ consists in encoding it in the form of a companion tuple $\bm{\tau}$ as in [35, 36]. Finite and discrete multidimensional spaces with more intricate configurations other than whole number multiplication in the form of $m_{1}\times m_{2}\times\dots\times m_{k}$, where $m,k\in\mathbb{N}$, can, for example, be encoded by node-unaligned companion tuples $\bm{\tau}_{ua}$ as in [36, Definition 2]. Nevertheless, the theoretical framework of our method is agnostic vis-a-vis any arbitrarily chosen computational scheme for encoding multidimensional spaces.
Each partition of a $n$-bit-length linear signal stream corresponds to a distinct configuration of dimension lengths, and therefore to a distinct multidimensional space in which the total additive dimension lengths remains upper bounded by the constant $n$. For example, a picture of 4$\times$4 pixels can be partitioned into other dimension-length configurations leading to different non-squared shapes, like 2$\times$8, 8$\times$2, or 1$\times$16. Once a row (or, alternatively, a column) is fixed in a bidimensional space, the remaining rows and columns are dependent on the initial row fixation. In other words, each bidimensional partition is a single-variable dependency, which is the main rationale underlying the sweep in our algorithms employed to achieve the results in Section 2.1.
More formally, in the case of a (finite and discrete) bidimensional space $\mathcal{S}_{2}$, one may have different combinations of values $\left|{\mathcal{S}_{2}}_{x}\right|$ and $\left|{\mathcal{S}_{2}}_{y}\right|$ for each partition. However, for any partition one has it that both $1\leq\left|{\mathcal{S}_{2}}_{x}\right|\leq n$ and $\lfloor\frac{n}{\left|{\mathcal{S}_{2}}_{x}\right|}\rfloor=\left|{\mathcal{S}_{2}}_{y}\right|$ hold. Thus, because each partition is single-variable dependent on ${\mathcal{S}_{2}}_{x}$, for any $n$-bit-length linear signal stream one can encode each of its $2$D-partitions in $\mathbf{O}\left(\log\left(n\right)\right)$ bits. The same applies to $3$D-partitions (or to any partition from a finite number of dimensions), except for an extra number of bits upper bounded by a partition-independent constant.
[…]
A (algorithmic) perturbation with low algorithmic complexity is one of these transformations where the algorithmic information content of the original object is mostly preserved under such perturbations. More formally, a low-complexity perturbation occurs when $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity, $x$ is the original encoded object, $y$ is the encoded object that results from the perturbation on $x$, and $\left|x\right|$ is the size of the object $x$.
Swapping all 0s for 1s, or swapping all black for white pixels in an image are examples of low-complexity perturbations (in these two cases in particular, one has it that $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(1\right)$). Randomly inserting or deleting a finite number of elements (such as creating or destroying edges in a graph or flipping a finite number of bits in a string) are also examples of low-complexity perturbations [26, 37, 24]. As we saw in the above paragraphs, changing the partition is also a low-complexity perturbation of the multidimensional space into which the object $x$ is embedded. This is because any new partition can always be encoded in $\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$ bits. Notice that a perturbation can change either the object (e.g., by flipping bits) or its multidimensional space (e.g., by reconfiguring the dimension lengths of a given partition to achieve another distinct partition) independently, or can change both at the same time. In any of these three options, except for the identity transformation, the encoded form of the message resulting from the perturbation will be distinct from the encoded form of the original message. Both changing the partition into which the object is embedded and introducing noise into the object itself (e.g., by flipping bits) are particular examples of (algorithmic) perturbations [26, 37, 24]. Notice that these properties of low-complexity perturbations underpin the methodology employed in Section 2.1.
$\mathbf{K}\left(\mathcal{S}\right)$ gives the algorithmic complexity of the arbitrarily chosen encoded form of the multidimensional space $\mathcal{S}$. $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)$ is the algorithmic complexity of the encoded object $y$ once the multidimensional space $\mathcal{S}$ is a priori known. $\mathbf{K}\left(y,\mathcal{S}\right)$ is the algorithmic complexity of the encoding of the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$. Note that there is no loss of generality in our proofs, nor any algorithmic information distortion (except for an independent constant), in employing here the usual notation $\mathbf{K}\left(w,z\right)$, which usually stands for the joint algorithmic complexity of a string $w$ and a string $z$. This is because there is an algorithm that can always encode the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$, if the shortest program that generates the encoding of the pair $\left(y,\mathcal{S}\right)$ is given as input to this algorithm; and there is another algorithm that can always return the encoded pair $\left(y,\mathcal{S}\right)$, if the shortest program that generates the encoded form of the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$ is given as input to this algorithm. For example, as shown in [35, 36], the composite edge set string of a multidimensional network is always computably retrievable from its characteristic string and the encoded form of its companion tuple; and both the characteristic string and the companion tuple of a multidimensional network are computably retrievable from the composite edge set string.
4.2 Toward a general theory for zero-knowledge one-way communication
Let $\mathbf{P}\left[\cdot,\cdot\right]$ be an arbitrarily chosen (computable) probability measure on the space of all possible (computably constructible) finite objects $y$ embedded into a multidimensional space $\mathcal{S}$. For example, $\mathbf{P}\left[y,\mathcal{S}\right]$ may be any probability measure value that is calculated for $y$ embedded into $\mathcal{S}$ by an arbitrarily chosen formal theory $\mathbf{F}$, encoding-decoding scheme, and decoding algorithm.
We know from the algorithmic coding theorem [38, 39, 40, 41] that
$\mathbf{K}\left(x\right)=$ (1)
$-\log\left(\mathbf{P}_{\mathbf{U}}\left[\text{``event }x\text{ occurs''}\right]\right)\pm\mathbf{O}(1)=$
$-\log\left(\mathbf{m}\left(x\right)\right)\pm\mathbf{O}(1)\text{ ,}$
holds, where: $\mathbf{m}\left(\cdot\right)$ is a maximal computably enumerable semimeasure; and $\mathbf{P}_{\mathbf{U}}\left[\cdot\right]$ denotes the universal a priori probability of an arbitrary event. $\mathbf{P}_{\mathbf{U}}$ can be understood as the probability of randomly generating (by an i.i.d. stochastic process) a prefix-free (or self-delimiting) program that outputs $x$. A computably enumerable semimeasure $\mathbf{m}\left(\cdot\right)$ is said to be maximal if, for any other computably enumerable semimeasure $\mu\left(\cdot\right)$ with domain defined for possible encoded objects, where $\sum\limits_{x\in\left\{0,1\right\}^{*}}\mu\left(x\right)\leq 1$, there is a constant $C>0$ (which does not depend on $x$) such that, for every encoded object $x$, $\mathbf{m}\left(x\right)\geq C\,\mu\left(x\right)\text{ .}$ Also note that the algorithmic coding theorem applies analogously to the conditional algorithmic complexity $\mathbf{K}(z\,|w)$.
We can explore the power and universality of the algorithmic coding theorem in order to investigate a lower bound for the universal (probability) distribution in comparison to an arbitrary probability measure $\mathbf{P}$. A chosen probability measure $\mathbf{P}$ may, or may not, assign a higher probability value to a correct multidimensional space. Thus, it is important to evaluate how much a universal approach such as the universal a priori probability can be wrong (i.e., diverges from $\mathbf{P}$) in the asymptotic limit. Indeed, the algorithmic coding theorem enables one to guarantee that the multiplicative error of
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]=\mathbf{P}_{\mathbf{U}}\left[y\middle|\mathcal{S}\right]$
with respect to
$\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]$
is bounded by the algorithmic probability of the inverse algorithmic perturbation $\mathcal{P}^{-1}$ itself. This means that the easier, or less complex, it is to retrieve $y$ embedded into $\mathcal{S}$ from $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$ the smaller the multiplicative error of employing $\mathbf{P}_{\mathbf{U}}$ instead of the arbitrarily chosen $\mathbf{P}$.
In the general case (where the multidimensional spaces may be much more complex than the objects themselves), even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended [35, 36]. (See also Section A.2 ) This poses a challenge to straightforward applications of the algorithmic information theory (AIT). In order to tackle such a challenge, we develop a method for multidimensional reconstruction of received low-dimensional signals in which one does not have access to knowledge about the encoding-decoding scheme or the original multidimensional space. Theorem 1 demonstrates the existence of sufficient conditions (or assumptions) that allow one to achieve an encoding/decoding-agnostic method for reconstructing the original multidimensional spaces. It gives rise to an agnostic method that does not need to assume that one a priori knows that the emitter agent’s multidimensional space resembles ours, is as complex as ours, or is isomorphic to the structures we currently deem fundamental in mathematics.
Theorem 1 (Theorem 11 in the Sup. Inf.).
Let $y$ be an arbitrary object embedded into an arbitrary multidimensional space $\mathcal{S}$ such that the following conditions
  • $\mathbf{K}\left(\mathcal{S}\right)=\mathbf{o}\big{(}\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\big{)}\text{ ,}$ (2)
  • $\mathbf{K}\left(\mathcal{P}\right)=\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}\text{ ,}$ (3)
  • and
  • $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\leq\mathbf{o}\big{(}\log\left(\left|y\right|\right)\big{)}$ (4)
hold, where $\mathcal{P}$ is any algorithmic perturbation that transforms $y$ embedded into $\mathcal{S}$ into another $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$. Let $\mathbf{P}^{\prime}\left[\cdot,\cdot\right]$ be the uniform probability measure over the space of all possible outcomes (in the form $\left(y^{\prime},\,\mathcal{S}^{\prime}\right)$) of algorithmic perturbations on $\left(y,\,\mathcal{S}\right)$ that satisfy Equation 3. Let $\mathbf{P}\left[\cdot,\cdot\right]$ be any arbitrarily chosen (computable) probability measure on the space of all possible (computably constructible) objects embedded into a multidimensional space. Then, the following strong asymptotic dominance
$\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]=\mathbf{o}\big{(}\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]\big{)}$ ” ] ) (5)
holds with probability (given by the uniform probability measure $\mathbf{P}^{\prime}$) arbitrarily close to $1$ as $\left|y\right|\to\infty$ for $\left(y^{\prime},\mathcal{S}^{\prime}\right)$ distinct from $\left(y,\mathcal{S}\right)$ satisfying Equation 3.
By applying AID to the problem of communicating messages so that one can perform a perturbation analysis on the received signal, Theorem 1 enables one to exploit the semi-computability of algorithmic complexity while taking advantage of the divergence brought about by the landscape of the approximate complexity values, landscape which results from the perturbations and from which we can find the best candidates to choose. Individually, the approximation to the complexity value of each object (and multidimensional space) is prone to the unknown error margins that follow from its semi-computability. However, because each pair object-space results from a perturbation, which is a particular case of algorithm, and each algorithm theoretically has an exact complexity value associated to it, one can retrieve a likely correct answer from a sufficiently large landscape of these perturbations.
As further discussed in Section A.5, we believe the results in this article not only contribute with a general method for interpreting messages in zero-knowledge one-way communication channels, but also establish fundamental connections (and posit open problems) with zero-knowledge proofs and the algorithmic counterparts of the maximum likelihood method and the noisy-channel coding theorem.
[…]
Acknowledgements
Felipe S. Abrahão acknowledges support from the São Paulo Research Foundation (FAPESP), grants $2021$/$14501$-$8$ and $2023$/$05593$-$1$.
Appendix A Supplementary Information
A.1 Introduction to the formal concepts
In what follows, we specify the scientific context and mathematical conditions that enable our results and claims.
For the present purposes, we say an object has “meaning” when the object conveyed in the received signal is actually embedded into (or grounded on) the multidimensional space (i.e., the context) that the emitter agent originally intended. Each (finite and discrete) multidimensional space $\mathcal{S}$—no matter how complex it is—can be univocally determined by the number of dimensions, the encoding of the set of elements for each dimension, and the ordering in which each dimension appears. For example: the unidimensional space $\mathcal{S}_{1}$ that takes values from the natural numbers can be univocally determined by only informing the length $\left|\mathcal{S}_{1}\right|$ in $\mathbf{O}\left(\log\left(\left|\mathcal{S}_{1}\right|\right)\right)$ bits; the bidimensional space $\mathcal{S}_{2}$ by a pair $\left(\left|{\mathcal{S}_{2}}_{x}\right|,\left|{\mathcal{S}_{2}}_{y}\right|\right)$ in $\mathbf{O}\left(\log\left(\left|{\mathcal{S}_{2}}_{x}\right|\right)\right)+\mathbf{O}\left(\log\left(\left|{\mathcal{S}_{2}}_{y}\right|\right)\right)+\mathbf{O}\left(1\right)$ bits, where ${\mathcal{S}_{2}}_{x}$ is the first dimension and ${\mathcal{S}_{2}}_{y}$ is the second dimension; and so on. In this manner, one is always able to (uniquely) decode the received message into sufficient information for completely determining the multidimensional space configuration. One example of encoding scheme for an arbitrary multidimensional space $\mathcal{S}$ is by encoding it in the form of a companion tuple $\bm{\tau}$ as in [35, 36]. Finite and discrete multidimensional spaces with more intricate configurations other than whole numbers multiplication in the form of $m_{1}\times m_{2}\times\dots\times m_{k}$, where $m,k\in\mathbb{N}$, can for example be encoded by node-unaligned companion tuples $\bm{\tau}_{ua}$ as in [36, Definition 2]. Nevertheless, as demonstrated in Section 2.2, the theoretical framework of our method is agnostic to any arbitrarily chosen computational scheme to encode multidimensional spaces.
Each partition of a $n$-bit-length linear signal stream corresponds to a distinct configuration of dimensions’ lengths, and therefore to a distinct multidimensional space in which the total additive dimensions’ lengths remains upper bounded by the constant $n$. For example, a picture of 4$\times$4 pixels can be partitioned into other dimension-length configurations leading to different non-squared shapes, like 2$\times$8, 8$\times$2, or 1$\times$16. Once a row (or, alternatively, a column) is fixed in a bidimensional space, the remaining rows and columns are dependent on the initial row fixation. In other words, each bidimensional partition is a single-variable dependency. More formally, in the case of a (finite and discrete) bidimensional space $\mathcal{S}_{2}$, one may have different combinations of values $\left|{\mathcal{S}_{2}}_{x}\right|$ and $\left|{\mathcal{S}_{2}}_{y}\right|$ for each partition. However, for any partition one has it that both $1\leq\left|{\mathcal{S}_{2}}_{x}\right|\leq n$ and $\lfloor\frac{n}{\left|{\mathcal{S}_{2}}_{x}\right|}\rfloor=\left|{\mathcal{S}_{2}}_{y}\right|$ hold. Thus, because each partition is single-variable dependent on ${\mathcal{S}_{2}}_{x}$, for any $n$-bit-length linear signal stream one can encode each of its $2$D-partitions in $\mathbf{O}\left(\log\left(n\right)\right)$ bits. The same applies to $3$D-partitions (or to any partition from a finite number of dimensions), except for an extra number of bits upper bounded by a partition-independent constant $\mathbf{O}\left(1\right)$, which basically corresponds to the algorithmic complexity of the minimal program that calculates the other dimensions given the respective single variable as input.
As formalised in [26] any transformation or perturbation of a finite (encoded) object into another finite (encoded) object is equivalent to an algorithmic transformation that takes the former as input and outputs the latter. (See also Section A.3.2). Thus, a (algorithmic) perturbation with low algorithmic complexity is one of these transformations whose algorithmic information content of the original object is mostly preserved under such perturbations. More formally, a low-complexity perturbation occurs when $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity, $x$ is the original encoded object, $y$ is the encoded object that results from the perturbation on $x$, and $\left|x\right|$ is the size of the object $x$.
Swapping all 0’s and 1’s, or swapping all black and white pixels in an image are examples of low-complexity perturbations (in these two cases in particular, one has it that $\mathbf{K}\left(y\middle|x\right)=\mathbf{O}\left(1\right)$). Randomly inserting or deleting a finite number of elements (such as creating or destroying edges in a graph or flipping a finite number of bits in a string) are also examples of low-complexity perturbations [26, 37, 24]. As we saw in the above paragraphs, changing the partition is also a low-complexity perturbation of the multidimensional space into which the object $x$ is embedded. This is because any new partition can always be encoded in $\mathbf{O}\left(\log\left(\left|x\right|\right)\right)$ bits. Thus, notice that a perturbation can change either the object (e.g., by flipping bits) or its multidimensional space (e.g., by reconfiguring the dimensions’ lengths of a given partition to achieve another distinct partition) independently, or can change both at the same time. In any of these three options, the encoded form of the message resulting from the perturbation will be distinct from the encoded form of the original message. Thus, notice that both changing the partition into which the object is embedded and introducing noise into the object itself (e.g., by flipping bits) are particular examples of (algorithmic) perturbations [26, 37, 24].
The theoretical framework and results presented in Section 2.2 demonstrate the method employed to achieve the empirical findings in Section A.4 is sound in the infinite asymptotic limit when computational resources are unbounded, and it can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Our empirical results corroborate the theoretical predictions of the theorems in Section 2.2, which in turn evinces that AID is capable of harnessing (under feasible computational resources) some of the algorithmic-informational properties that the following theorems demonstrate in ideal and generalised scenarios. As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, and most prominently for high algorithmically probable objects.
In Section A.2, we present the theoretical limitations in the general case, which in turn highlights the sufficient conditions that should be satisfied in order to avoid these limitations as we will investigate in the theory developed in Section 2.2.
A.2 Avoiding algorithmic information distortions in arbitrarily complex multidimensional spaces
In [35, 36] it is investigated how much of the irreducible information content of a multidimensional network and its isomorphic monoplex network (i.e., its isomorphic graph) is preserved during such an isomorphism transformation. In fact, isomorphisms are demonstrated to not preserve algorithmic information in the general case. The algorithmic information of a multidimensional network might be exponentially distorted with respect to the algorithmic information of its isomorphic (monoplex) network (see [36, Theorem 12]). We refer to an algorithmic information distortion when the algorithmic information content of an object (e.g., measured by the prefix algorithmic complexity of an encoded form of the object) in a certain context (e.g., the multidimensional space the object is embedded into) sufficiently differs from the algorithmic information content of the same object in a distinct context. In this regard, algorithmic complexity oscillations may arise from a series of distortions produced by a large enough collection of distinct contexts into which the same object is embedded.
From the well-known invariance in algorithmic information theory (AIT) [38, 39, 40, 41], one for example knows that every positive or negative oscillation resulting from changing the computation model (or universal programming language) is pairwise upper and lower bounded by a constant that does not depend on the object output by this pair of machines. Another well-known example of (plain) algorithmic complexity oscillations are those present in initial segments of infinite strings [40].
As shown in Section A.4, other examples are the oscillations measured by the plots presented in the figures, particularly those image plots for which one has the $x$-axis representing the distinct partitions we estimated the algorithmic complexity of the messages into, respectively. Notice that the downward spikes one can see in Section A.4 are examples of algorithmic complexity (negative) oscillations such that for the decoding problem we are interested in, they indicate the partition that corresponds to the correct multidimensional space the message intended.
The results in [35, 36] demonstrate a fundamental limitation that mathematical and computational methods should be aware of in order to properly reconstruct the original multidimensional space from a receiver. To demonstrate this limitation, take for example received signals that encode (either in $1$D or $2$D signal streams) adjacency matrices of the (monoplex) networks whose isomorphic counterparts are multidimensional. For any arbitrarily chosen formal theory and decoding algorithm, the theorems in [35, 36] imply that there is a sufficiently large multidimensional network whose multidimensional space cannot be retrieved from the adjacency matrix of the network by the respective formal theory and/or decoding algorithm. This occurs because the algorithmic information distortions (due to being exponential) simply outgrow any prediction or output capability of the formal theory together with the decoding algorithm, should the network be given as input [36, Theorem 12].
Such distortions demonstrate that in the general case where the multidimensional spaces may be much more complex than the objects themselves, there is a fundamental limitation for any formal theory with regard to the original multidimensional space to which a received message containing an object embedded into a very simple multidimensional space (e.g., a linear one, such as a signal stream) in fact corresponds [42]. This implies that even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended. In addition to the distinctive feature of AID in comparison to classical AIT that we will discuss in Section A.5.1, these algorithmic distortions also pose a challenge to straightforward applications of the algorithmic coding theorem [38, 39, 40, 41], universal (Solomonoff) induction, or minimum description length methods [27, 40]. This challenge may be even more aggravated in the case of arbitrarily large signal streams which correspond to partial messages made available to the receiver agent so that the respective intended messages will be only completed in larger signal streams [43].
Intuitively, it is straightforward for us humans to conceive and construct an object, and then to transmit a signal with this object encoded in the message the signal is conveying, so that this object is much more complex than the space-time $4$-dimensional space that we are used to. However, in the general case either the multidimensional space of a unknown source can be much more complex than humans are used to or the object that the unknown emitter is trying to convey can be relatively simple in comparison to this yet unknown multidimensional space. Both cases can be affected by the limitations brought about by the aforementioned algorithmic information distortions.
Therefore, in order to circumvent such theoretical limitations, one needs to investigate additional conditions to be met so one can correctly infer the message given that the receiver has no priori knowledge about the source. In other words, with the purpose of developing methods for multidimensional reconstruction of received low-dimensional signals in which one does not have access to knowledge about the encoding-decoding scheme or the original multidimensional space, one should look for conditions or assumptions that avoid such distortions. One most obvious approach to avoiding them is to limit the algorithmic complexity of the multidimensional space itself with respect to the algorithmic complexity of the object. As we will demonstrate in Section 2.2, a proper formalisation of this intuition is indeed sufficient for achieving an encoding/decoding-agnostic method for reconstructing the original multidimensional spaces. Thus, the proven existence of these sufficient conditions in turn highlight the generality and contributions of our present results. They present an agnostic method that does not need to assume that one a priori knows that the emitter agent’s multidimensional space resembles ours, is as complex as ours, or is isomorphic to the structures we currently deem fundamental in mathematics.
A.3 Theoretical results
A.3.1 Algorithmic information of encoded objects and multidimensional spaces
Let $\mathbf{K}\left(\mathcal{S}\right)$ denote the algorithmic complexity of the arbitrarily chosen encoded form of the multidimensional space $\mathcal{S}$; $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)$ denote the algorithmic complexity of the encoded object $y$ once the multidimensional space $\mathcal{S}$ is a priori known; and $\mathbf{K}\left(y,\mathcal{S}\right)$ denote the algorithmic complexity of the encoding of the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$. Note that there is no loss of generality in our following proofs, nor any algorithmic information distortion (except for an independent constant), in employing here the usual notation $\mathbf{K}\left(w,z\right)$, which usually stands for the joint algorithmic complexity of a string $w$ and a string $z$. This is because there is an algorithm that can always encode the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$, if the shortest program that generates the encoding of the pair $\left(y,\mathcal{S}\right)$ is given as input to this algorithm; and there is another algorithm that can always return the encoded pair $\left(y,\mathcal{S}\right)$, if the shortest program that generates the encoded form of the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$ is given as input to this algorithm. For example, as shown in [35, 36], the composite edge set string of a multidimensional network is always computably retrievable from its characteristic string and the encoded form of its companion tuple; and both the characteristic string and the companion tuple of a multidimensional network are computably retrievable from the composite edge set string.
As usual, let $f(x)=\mathbf{O}(g(x))$, $f(x)=\mathbf{o}(g(x))$, $f(x)=\bm{\omega}(g(x))$, and $f(x)\sim g(x)$ denote asymptotic dominances.
The following Lemma 2 demonstrates a sufficient condition that assures that any algorithmic information distortion—which might occur due to changing the multidimensional space into which the object should be embedded—vanishes in the limit as the object itself becomes sufficiently complex with respect to the multidimensional space.
Lemma 2.
Let $\mathcal{S}$ be an arbitrary multidimensional space. Let $y$ be an arbitrary object embedded into $\mathcal{S}$. Therefore, if
$\mathbf{K}\left(\mathcal{S}\right)=\mathbf{o}\big{(}\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\big{)}$ (6)
holds, then for any $y$ with sufficiently large $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)$, we have it that
$\mathbf{K}\left(y,\,\mathcal{S}\right)\sim\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\text{ .}$ (7)
i.e., the algorithmic complexity of the object $y$ univocally embedded into the multidimensional space $\mathcal{S}$ is of the same order of the algorithmic complexity of the object in the given multidimensional space.
Proof.
From basic properties of $\mathbf{K}\left(\cdot\right)$ [38, 39, 40, 41], we have that $\mathbf{K}(w\,|z)\leq\mathbf{K}(w)+\mathbf{O}(1)\leq\mathbf{K}(w,z)+\mathbf{O}(1)\leq\;\mathbf{K}(z)+\mathbf{K}(w\,|z)+\mathbf{O}(1)$ holds for any $w$ and $z$. Finally, from this latter inequality and the condition $\mathbf{K}\left(\mathcal{S}\right)=\mathbf{o}\left(\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\right)$ in Equation (6), we will have it that
$\lim\limits_{\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\to\infty}\frac{\mathbf{K}(y,\mathcal{S})}{\mathbf{K}(y\,|\mathcal{S})}=1\text{ .}$
By logarithmically compressible objects we refer to those objects whose encoded form can be compressed up to a logarithmic term (i.e., $\mathbf{K}\left(x\right)\leq\mathbf{O}\big{(}\log\left(\left|x\right|\right)\big{)}$, where $x$ is an encoded object) as a function of the size of the object. For the purposes of the present paper, $\left|x\right|$ is the length of the linear signal stream.
Corollary 3 then straightforwardly shows that, in case the conditions of Lemma 2 hold, the logarithmic compressibility of the message is preserved:
Corollary 3.
Let $y$ and $\mathcal{S}$ satisfy Lemma 2 such that the length of the linear signal stream $\left|y\right|$ is sufficiently large and
$\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\leq\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}\text{ .}$ (8)
Then,
$\mathbf{K}\left(y,\,\mathcal{S}\right)=\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}$ (9)
holds, i.e., the algorithmic complexity of the object univocally embedded into the multidimensional space remains logarithmically compressible.
We shall see in Corollary 6 (following from Lemma 4) that these conditions in Corollary 3 also guarantee that the algorithmic complexities indeed remain invariant to low-algorithmic-complexity perturbations. In addition, notice that in our final Theorem 11 we will employ the stronger version of the asymptotic dominance in the form $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\leq\mathbf{o}\left(\log\left(\left|y\right|\right)\right)$, which in turn also satisfies the condition of the above Corollary 3 in Equation 8.
A.3.2 Robustness of non-random messages to low-complexity perturbations
[…]
Lemma 4.
Let $y$ and $\mathcal{S}$ satisfy Lemma 2 such that $\left|y\right|$ is sufficiently large. Then,
$\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\sim$ (10)
$\mathbf{K}\left(y,\,\mathcal{S}\right)\leq$
$\mathbf{K}\Big{(}\mathbf{U}\big{(}\left<\left<y,\,\mathcal{S}\right>,\mathcal{P}\right>\big{)}\Big{)}+\mathbf{K}\left(\mathcal{P}^{-1}\right)+\mathbf{O}(1)=$
$\mathbf{K}\Big{(}y^{\prime},\mathcal{S}^{\prime}\Big{)}+\mathbf{K}\left(\mathcal{P}^{-1}\right)+\mathbf{O}(1)$
and
$\mathbf{K}\Big{(}y^{\prime},\mathcal{S}^{\prime}\Big{)}=$ (11)
[…]
$\mathbf{K}\left(y,\,\mathcal{S}\right)+\mathbf{K}\left(\mathcal{P}\right)+\mathbf{O}(1)\sim$
$\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)+\mathbf{K}\left(\mathcal{P}\right)+\mathbf{O}(1)$
hold, where: $\mathcal{P}$ is an algorithmic perturbation that transforms $y$ embedded into $\mathcal{S}$ into another arbitrary $y^{\prime}$ embedded into another arbitrary $\mathcal{S}^{\prime}$; and $\mathcal{P}^{-1}$ is the respectively inverse algorithmic perturbation that transforms $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$ back into $y$ embedded into $\mathcal{S}$.
Proof.
First, note that since $y$, $y^{\prime}$, $\mathcal{S}$, $\mathcal{S}^{\prime}$ are both finite and encodable objects by assumption, then both the existence of $\mathcal{P}$ and $\mathcal{P}^{-1}$ is guaranteed. Then, the desired proof of both Equations 10 and 11 follows directly from Lemma 2 and the basic property [40, 41, 38] of $\mathbf{K}\left(\cdot\right)$ which demonstrates that $\mathbf{K}\left(f(w)\right)\leq\mathbf{K}\left(w\right)+\mathbf{K}\left(\#_{f}\right)$ holds for any $w$ and $f$, where $f$ is a computable function and #$\#_{f}$ denotes its encoding in $\mathbf{U}$. ∎
Corollary 5 then establishes that an algorithmic perturbation can (in the worst case scenario of high-complexity perturbations) increase the asymptotic dominance class of the algorithmic complexity of the original message. This is congruent with the forthcoming Corollary 10, which assumes the condition of $\mathbf{K}\left(y^{\prime},\mathcal{S}^{\prime}\right)$ being sufficiently larger than $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)$ in order to guarantee a strong asymptotic dominance.
Corollary 5.
Let $y$ and $\mathcal{S}$ satisfy Corollary 3 and Lemma 4 such that $\left|y\right|$ is sufficiently large. Then,
$\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)-\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}-\mathbf{O}(1)\leq\mathbf{K}\Big{(}y^{\prime},\mathcal{S}^{\prime}\Big{)}$ (12)
holds, where the algorithmic perturbation $\mathcal{P}$ transforms $y$ embedded into $\mathcal{S}$ into another $y^{\prime}$ embedded into another $\mathcal{S}^{\prime}$.
In addition, in the case of low-complexity perturbations, the asymptotic dominance class of the algorithmic complexity of the perturbed—and, therefore, possibly grounded on a slightly wrong multidimensional space—message remains as logarithmically compressible as the correct one:
Corollary 6.
Let $y$ and $\mathcal{S}$ satisfy Corollary 5 such that $\left|y\right|$ is sufficiently large and $\mathbf{K}\left(\mathcal{P}\right)=\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}$. Then,
$\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)-\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}-\mathbf{O}(1)\leq$ (13)
$\mathbf{K}\Big{(}y^{\prime},\mathcal{S}^{\prime}\Big{)}=$
$\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}$
holds, where the algorithmic perturbation $\mathcal{P}$ transforms $y$ embedded into $\mathcal{S}$ into another $y^{\prime}$ embedded into another $\mathcal{S}^{\prime}$.
A.3.3 Robustness to arbitrary probability measures and to perturbations of objects and multidimensional spaces
Let $\mathbf{P}\left[\cdot,\cdot\right]$ be an arbitrarily chosen (computable) probability measure on the space of all possible (computably constructible) finite objects $y$ embedded into a multidimensional space $\mathcal{S}$. For example, $\mathbf{P}\left[y,\mathcal{S}\right]$ may be any probability measure value that is calculated for $y$ embedded into $\mathcal{S}$ by an arbitrarily chosen formal theory $\mathbf{F}$, encoding-decoding scheme, and decoding algorithm.
We know from the algorithmic coding theorem [38, 39, 40, 41] that
$\mathbf{K}\left(x\right)=$ (14)
$-\log\left(\mathbf{P}_{\mathbf{U}}\left[\text{``event }x\text{ occurs''}\right]\right)\pm\mathbf{O}(1)=$
$-\log\left(\mathbf{m}\left(x\right)\right)\pm\mathbf{O}(1)\text{ ,}$
holds, where: $\mathbf{m}\left(\cdot\right)$ is a maximal computably enumerable semimeasure; and $\mathbf{P}_{\mathbf{U}}\left[\cdot\right]$ denotes the universal a priori probability of an arbitrary event. $\mathbf{P}_{\mathbf{U}}$ can be understood as the probability of randomly generating (by an i.i.d. stochastic process) a prefix-free (or self-delimiting) program that outputs $x$. 222 In case the prefix-free (or self-delimiting) programming language constitutes a semimeasure instead of measure, one can straightforwardly apply a normalizing constant as in [44, Definition 3.6, Section 3.2.1]. A computably enumerable semimeasure $\mathbf{m}\left(\cdot\right)$ is said to be maximal if, for any other computably enumerable semimeasure $\mu\left(\cdot\right)$ with domain defined for possible encoded objects, where $\sum\limits_{x\in\left\{0,1\right\}^{*}}\mu\left(x\right)\leq 1$, there is a constant $C>0$ (which does not depend on $x$) such that, for every encoded object $x$, $\mathbf{m}\left(x\right)\geq C\,\mu\left(x\right)\text{ .}$ Also note that the algorithmic coding theorem applies analogously to the conditional algorithmic complexity $\mathbf{K}(z\,|w)$.
We can explore the power and universality of the algorithmic coding theorem in order to investigate a lower bound for the universal (probability) distribution in comparison to an arbitrary probability measure $\mathbf{P}$. A chosen probability measure $\mathbf{P}$ may, or may not, assign a higher probability value to a correct multidimensional space. Thus, it is important to evaluate how much a universal approach such as the universal a priori probability can be wrong (i.e., diverges from $\mathbf{P}$) in the asymptotic limit. Indeed, the algorithmic coding theorem together with Lemma 4 enables the following Theorem 7 to guarantee that the multiplicative error of
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]=\mathbf{P}_{\mathbf{U}}\left[y\middle|\mathcal{S}\right]$
with respect to
$\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]$
is bounded by the algorithmic probability of the inverse algorithmic perturbation $\mathcal{P}^{-1}$ itself. This means that the easier, or less complex, it is to retrieve $y$ embedded into $\mathcal{S}$ from $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$ the smaller the multiplicative error of employing $\mathbf{P}_{\mathbf{U}}$ instead of the arbitrarily chosen $\mathbf{P}$.
Theorem 7.
Let $y^{\prime}$ be an arbitrary object and $\mathcal{S}^{\prime}$ be an arbitrary multidimensional space. Let an arbitrary $y$ embedded into arbitrary $\mathcal{S}$ satisfy Lemma 4. Then,
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]\geq$ ” ] ≥ (15)
$\frac{1}{2^{\mathbf{K}\left(y,\,\mathcal{S}\right)\pm\mathbf{O}(1)}}\geq$
$\frac{1}{2^{\mathbf{K}\left(y^{\prime},\,\mathcal{S}^{\prime}\right)+\mathbf{K}\left(\mathcal{P}^{-1}\right)\pm\mathbf{O}(1)}}\geq$
$\frac{\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]}{2^{\mathbf{K}\left(\mathcal{P}^{-1}\right)+\mathbf{O}(1)}}$
holds, where $\mathcal{P}^{-1}$ is the inverse algorithmic perturbation that transforms $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$ back into $y$ embedded into $\mathcal{S}$.
Proof.
First, note that the inequality
$\frac{1}{2^{\mathbf{K}\left(y,\,\mathcal{S}\right)\pm\mathbf{O}(1)}}\geq\frac{1}{2^{\mathbf{K}\left(y^{\prime},\,\mathcal{S}^{\prime}\right)+\mathbf{K}\left(\mathcal{P}^{-1}\right)\pm\mathbf{O}(1)}}$ (16)
follows directly from Lemma 4. Secondly, we will have from the algorithmic coding theorem (see Equation (14))333 Actually, for this step in particular, one just needs to apply the definition of the universal a priori probability $\mathbf{P}_{\mathbf{U}}$ to demonstrate that it dominates the algorithmic probability $2^{-\mathbf{K}\left(\cdot\right)}$ from above. and Lemma 2 that
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]=$ ” ] = (17)
$\mathbf{P}_{\mathbf{U}}\left[y\middle|\mathcal{S}\right]\geq$
$\frac{1}{2^{\mathbf{K}\left(y\,\middle|\mathcal{S}\right)+\mathbf{O}(1)}}\sim$
$\frac{1}{2^{\mathbf{K}\left(y,\,\mathcal{S}\right)+\mathbf{O}(1)}}\text{ .}$
Finally, from the algorithmic coding theorem, we will have it that
$\frac{1}{2^{\mathbf{K}\left(y^{\prime},\,\mathcal{S}^{\prime}\right)+\mathbf{O}(1)}}\geq$ (18)
$\frac{1}{2^{\mathbf{O}(1)}}\,\mathbf{P}_{\mathbf{U}}\left[y^{\prime},\,\mathcal{S}^{\prime}\right]\geq$
$\frac{1}{2^{\mathbf{O}(1)}}\,\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]\text{ .}$
Now, directly following from Theorem 7, one demonstrates Corollary 8, which shows that one can only wrongly assign a probability for $y$ up to $2^{\mathbf{o}\left(\mathbf{K}\left(y\right)\right)+\mathbf{O}(1)}$ less likely than $\mathbf{P}$ can assign (except for the algorithmic probability of reverting the perturbation that might have introduced noise into $y$, perturbation which resulted in $y^{\prime}$). This is a multiplicative term that does not depend on $\mathcal{S}$ nor $\mathcal{S}^{\prime}$. And such a lower bound holds whether $\mathcal{S}=\mathcal{S}^{\prime}$ or not, i.e., whether $\mathcal{S}^{\prime}$ is the correct multidimensional space $\mathcal{S}$, or any other wrong one.
Corollary 8.
Let $y^{\prime}$ be an arbitrary object and $\mathcal{S}^{\prime}$ be an arbitrary multidimensional space. Let an arbitrary $y$ embedded into arbitrary $\mathcal{S}$ satisfy Lemma 4. Then,
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]\geq$ ” ] ≥ (19)
$\frac{1}{2^{\mathbf{K}\left(y,\,\mathcal{S}\right)\pm\mathbf{O}(1)}}\geq$
$\frac{1}{2^{\mathbf{K}\left(y^{\prime},\,\mathcal{S}^{\prime}\right)+\mathbf{K}\left(y\middle|y^{\prime},\mathcal{S}\right)+\mathbf{o}\left(\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\right)\pm\mathbf{O}(1)}}\geq$
$\frac{\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]}{2^{\mathbf{K}\left(y\middle|y^{\prime},\mathcal{S}\right)+\mathbf{o}\left(\mathbf{K}\left(y\middle|\mathcal{S}\right)\right)+\mathbf{O}(1)}}$
holds.
That is, Corollary 8 says that the multiplicative error of
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]$ ” ]
with respect to
$\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]$
increases asymptotically slower than the algorithmic probability of $y$, except for the algorithmic probability of generating $y$ from $y^{\prime}$ in the given $\mathcal{S}$, where $y^{\prime}$ may e.g. result from noise introduced into $y$.
Hence, in the case without any noise together with a logarithmic compressibility (from Corollary 5), one can directly show in Corollary 9 that the assumption of logarithmic compressibility ensures that searching for the multidimensional space that minimises the algorithmic complexity can only produce errors that vanishes linearly in the asymptotic limit in comparison to the algorithmic complexity of the object itself. This holds because in this case the multiplicative error of employing the universal distribution instead of any other arbitrarily chosen probability measure (which assigns higher likelihood to the multidimensional space $\mathcal{S}^{\prime}$) becomes strongly dominated by the algorithmic probability of the object $y$ itself, which in turn decreases linearly with the size of $y$.
Corollary 9.
Let $\mathcal{S}^{\prime}$ be an arbitrary multidimensional space. Let an arbitrary $y$ embedded into arbitrary $\mathcal{S}$ satisfy Corollary 5. Then, there is a constant $C$ such that the multiplicative error of
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]$ ” ]
with respect to
$\mathbf{P}\left[y,\mathcal{S}^{\prime}\right]$
increases asymptotically slower than
$\frac{C}{\left|y\right|}\text{ .}$
A.3.4 Bias toward low complexity in zero-knowledge one-way communication
So far, we have investigated the worst-case scenarios in which $\mathbf{P}_{\mathbf{U}}$ may be underestimating the probability measure $\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]$ of your choice up to a multiplicative term lower bounded by $2^{-\mathbf{K}\left(\mathcal{P}^{-1}\right)-\mathbf{O}(1)}$. However, as indicated by our empirical results, we can demonstrate in Corollary 10, which in turn follows directly from Corollary 6 given that the conditions of Corollary 8 (or, equivalently, Theorem 7) are also satisfied, that the original multidimensional space with the lowest algorithmic complexity can be set apart from the wrong ones, should the wrong ones sufficiently move (i.e., transform or perturb) the object and/or the multidimensional space toward randomness in comparison to $\mathbf{K}\left(y\middle|\mathcal{S}\right)$. In addition, from the results demonstrated in the previous Section A.3.3, note that this phenomenon holds for any arbitrarily chosen computable probability measure $\mathbf{P}$ as long as the algorithmic complexity of the object $y$ is sufficiently large.
Corollary 10.
Let an arbitrary $y$ embedded into arbitrary $\mathcal{S}$ satisfy Corollary 6. Let $y^{\prime}$ be an arbitrary object and $\mathcal{S}^{\prime}$ be an arbitrary multidimensional space such that
$\mathbf{K}\left(y^{\prime},\mathcal{S}^{\prime}\right)=\bm{\omega}\big{(}\mathbf{K}\left(y\middle|\mathcal{S}\right)\big{)}\text{ .}$ (20)
Then,
$\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]=\mathbf{o}\big{(}\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]\big{)}$ ” ] ) (21)
holds.
In summary, Corollary 10 tells us that decoding the message in the original multidimensional space $\mathcal{S}$ chosen from the one in which the message has the lowest algorithmic complexity (i.e., highest algorithmic probability, as assured by the universal distribution defined by $\mathbf{P}_{\mathbf{U}}$) becomes eventually more probable to be correct than decoding the message in an distinct multidimensional space $\mathcal{S}^{\prime}$, with or without noise. The key idea of Corollary 10 is that the larger the algorithmic complexity of the closer-to-randomness message candidate in the wrong multidimensional space with respect to the algorithmic complexity of the message in the correct multidimensional space, the larger the probability of the correct multidimensional space with respect to the probability of the wrong closer-to-randomness multidimensional space. Thus, finding the multidimensional space that minimises the algorithmic complexity the most, eventually increases the probability of finding the correct one in comparison to the probability of finding the wrong closer-to-randomness one. Notice that this result in Corollary 10 holds even if one assumes an arbitrarily different probability measure as starting point. In addition, the strong asymptotic dominance renders any constant that may appear in the algorithmic coding theorem irrelevant as $\left|y\right|\to\infty$.
The condition for this to hold is that the algorithmic complexity of the object (with or without noise) in the wrong multidimensional space should increase (i.e., move the message toward algorithmic randomness) sufficiently fast in comparison to the algorithmic complexity of the message in a correct multidimensional space. In addition to the assumption (see Lemma 2) of relatively low algorithmic complexity of the multidimensional space with respect to that of the object, one can assume a message should contain a large amount of redundancies, and thus should be highly compressible with respect to its linear signal stream length $\left|y\right|$, such as the case in which $\mathbf{K}\left(y\middle|\mathcal{S}\right)=\mathbf{O}\left(\log\log\left(\left|y\right|\right)\right)$ holds (which also satisfies the condition in Corollary 3). In this manner, most low-complexity perturbations that satisfy the condition in Corollary 6 will transform the object and/or partition by sufficiently moving it toward algorithmic randomness so that the conditions of the Corollary 10 become satisfied. This is because we know from AIT that the number of low-complexity perturbations for which $\mathbf{K}\left(y^{\prime},\mathcal{S}^{\prime}\right)=\bm{\omega}\big{(}\log\log\left(\left|y\right|\right)\big{)}$ holds strongly dominates the number of those for which one has it that $\mathbf{K}\left(y^{\prime},\mathcal{S}^{\prime}\right)=\mathbf{O}\left(\log\log\left(\left|y\right|\right)\right)$, which constitutes one of the main ideas behind the following Theorem 11.
Finally, one can then demonstrate and formalise the previous paragraph in its general form in our last Theorem 11, which encompasses the previous conditions present in the previous lemmas, theorems, and corollaries so as to achieve a proof that underpins both our empirical method and theoretical framework employed in the present work, proving that these are assured to work in the asymptotic limit optimal case of algorithmic complexity:
Theorem 11.
Let $y$ be an arbitrary object embedded into an arbitrary multidimensional space $\mathcal{S}$ such that the following conditions
  • $\mathbf{K}\left(\mathcal{S}\right)=\mathbf{o}\big{(}\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\big{)}\text{ ,}$ (22)
  • $\mathbf{K}\left(\mathcal{P}\right)=\mathbf{O}\big{(}\log\left(\left|y\right|\right)\big{)}\text{ ,}$ (23)
  • and
  • $\mathbf{K}\left(y\,\middle|\,\mathcal{S}\right)\leq\mathbf{o}\big{(}\log\left(\left|y\right|\right)\big{)}$ (24)
hold, where $\mathcal{P}$ is any algorithmic perturbation that transforms $y$ embedded into $\mathcal{S}$ into another $y^{\prime}$ embedded into $\mathcal{S}^{\prime}$. Let $\mathbf{P}^{\prime}\left[\cdot,\cdot\right]$ be the uniform probability measure over the space of all possible outcomes (in the form $\left(y^{\prime},\,\mathcal{S}^{\prime}\right)$) of algorithmic perturbations on $\left(y,\,\mathcal{S}\right)$ that satisfy Equation 23. Let $\mathbf{P}\left[\cdot,\cdot\right]$ be any arbitrarily chosen (computable) probability measure on the space of all possible (computably constructible) objects embedded into a multidimensional space. Then, the following strong asymptotic dominance
$\mathbf{P}\left[y^{\prime},\mathcal{S}^{\prime}\right]=\mathbf{o}\big{(}\mathbf{P}_{\mathbf{U}}\left[\text{``event }y\text{ occurs in a given }\mathcal{S}\,\text{''}\right]\big{)}$ ” ] ) (25)
holds with probability (given by the uniform probability measure $\mathbf{P}^{\prime}$) arbitrarily close to $1$ as $\left|y\right|\to\infty$ for $\left(y^{\prime},\mathcal{S}^{\prime}\right)$ distinct from $\left(y,\mathcal{S}\right)$ satisfying Equation 23.
Proof.
From Equations (22) and (24) and the basic properties of $\mathbf{K}\left(\cdot\right)$, one knows that there are only up to
$\mathbf{o}\left(\left|y\right|\right)$ (26)
possible distinct algorithmic perturbations $\mathcal{P}$ for which
$\mathbf{K}\Big{(}y^{\prime},\mathcal{S}^{\prime}\,\bigl{|}\,y,\,\mathcal{S}\Big{)}=\mathbf{K}\Big{(}\mathbf{U}\big{(}\left<\left<y,\,\mathcal{S}\right>,\mathcal{P}\right>\big{)}\,\bigl{|}\,y,\,\mathcal{S}\Big{)}\leq\mathbf{O}\big{(}\mathbf{K}\left(y,\,\mathcal{S}\right)\big{)}$ (27)
holds.444 Notice that the basic inequality $\mathbf{K}(w,z)\leq\;\mathbf{K}(z)+\mathbf{K}(w\,|z)+\mathbf{O}(1)$ employed here was the same employed in the proof of Lemma 2. From Equation 23 and the properties of algorithmic complexity in AIT, one also knows that there are at least
$\frac{\left|y\right|}{2^{\mathbf{O}(1)}}$ (28)
possible distinct outcomes $\left(y^{\prime},\,\mathcal{S}^{\prime}\right)$ for which Equation 20 holds. Then, we will have it that
[…]
as $\left|y\right|\to\infty$, which is a convergence that holds because of Equations 26 and 28 and $\mathbf{P}^{\prime}$ being a uniform probability measure by assumption. Now, notice that Equation 29 means that, Equation 20 in Corollary 10 is satisfied with probability $\mathbf{P}^{\prime}$ arbitrarily close to $1$ as $\left|y\right|\to\infty$. Finally, since all the other conditions for Corollary 10 are satisfied because the stated conditions in Theorem 11 directly satisfy those in the lemmas, theorems, and corollaries on which Corollary 10 depends, we achieve the desired proof that Equation 25 holds with probability $\mathbf{P}^{\prime}$ arbitrarily close to $1$ as $\left|y\right|\to\infty$. ∎
In other words, Theorem 11 tells us that once certain sufficient conditions for the compressibility of objects, multidimensional spaces, and perturbations (e.g., noise) are met, one-way communication to the receiver becomes possible even without prior knowledge about the decoding scheme chosen by the emitter. Being able to calculate the optimal values of algorithmic complexity in this case implies that one can first simply look into the landscape of all possible partitions and low-complexity noise the receiver has access to while assuming no bias toward particular objects and partitions in this landscape. Then, one can infer with probability as high as one wishes that the message with the lowest algorithmic complexity is arbitrarily more likely to be correct than any other arbitrary computable method to assign probabilities to message candidates.
A.4 Empirical results
A.4.1 Perturbation analysis of compressed data
To demonstrate the relationship between compression, algorithmic complexity, and encoding methods, we took real sentences from Darwin’s “Origin of Species” and binarised them in two different ways according to arbitrary ASCII-to-binary functions. In one encoding (left column in Fig. 2), ASCII vowels are assigned a 1 and all others 0. For the other encoding (right column), spaces are assigned a 1 and the rest 0. The compression, block entropy, and block decomposition method (BDM) are measured for each string. Then, for each string, each value is perturbed (1 to a 0, or vice-versa) one at a time. The resulting change in values from the original binary sequence to the perturbed ones are shown as different coloured lines for each panel.
The results in Fig. 2 demonstrate that most single-bit perturbations in a sequence that encodes a message result in a more random-looking sequence. These results vary depending on what mapping we choose to binarise the sequences, but in general, any perturbation generally increases randomness across various mappings (we also explored other arbitrary binarisation functions, with similar results). In other words, this example shows that sequences that carry meaningful content are quantitatively different according to various measures. Any deviation (i.e., perturbation) from the encoded sequences causes the randomness to increase, which suggests that an incoming signal with low randomness is more likely to encode a message than one that has more randomness.
Figure 2: Algorithmic perturbation analysis on short sentences of only 400 characters from Darwin’s “Origin of Species” are converted from ASCII into their corresponding binary values. Changing a single character with probability $1/n$ for growing $n$ from $n=1$ (original test, axis $x=0$) to $n=4$ (a quarter of the text is mapped to a random character from the same vocabulary of the original text, axis $x=100$ letters changed or perturbed) causes the resulting size of the compressed sequence index to increase, with BDM the most sensitive and Shannon entropy unable to capture any difference. This is because signals carrying meaning are far removed from randomness, and random perturbations make the text more random [37] even when the methods know nothing about words, grammar or anything linguistic.
Fig. 3 demonstrates the same phenomenon shown in non-random text but in audio signals, the results using an audio recording of the words spoken from Apollo 13, “Houston, we’ve had a problem” transmitted on April 13, 1970 at a sample rate frequency of 11025 Hz on a single channel. When compressed, the message is 106 250 bytes and 106 320 characters long. Here, the message was scrambled several times. The histograms on the bottom left show the different compression lengths in bytes and string lengths of the scrambled messages. Each of these scramblings results in an increase in randomness. Additionally, if only the beginning and end of each word are perturbed, the resulting randomness increases smoothly with each perturbation, as shown in the bottom right plot. These results indicate that perturbations of the original message, including scrambling elements of the original message in a different order, will result in a message of higher complexity and thus greater randomness. As a result, the original encoding of the message (the one with an interpretable meaning) is the one with the lowest complexity.
Figure 3: Top: Sound waves of the words spoken from Apollo 13, “Houston, we’ve had a problem.” transmitted on April 13, 1970 at a sample rate frequency of 11025 Hz on a single channel. Bottom left: Histograms of the change in compression lengths of scrambled versions of the same message from the length of the original compressed message. Bottom right: Small perturbations of the original message by scrambling only the beginning and end of each word shows a smooth transition from low to high randomness. The lowest complexity signal indicates the correct (original) signal. The file was processed in FLAC format (Free Lossless Audio Codec) from a lossless file, with no audio data discarded.
A.4.2 On the information content of the Arecibo message
In 1974, the bitmap image in Fig. 4 was sent into space as a radio signal from the Arecibo radio telescope. At the left-hand end of the image is a version of the pattern of digits from page 117– but it is distorted so it has no obvious nested structure. In the image, there are atomic numbers for various elements and bitvectors for components of DNA. Under these, there are rough pictorial representations of a DNA molecule, a human, and the telescope. All these images seem to depend almost completely on human visual encoding/decoding conventions. Without any sort of human context, including any indication that these are pictorial representations based on human vision, their meaning would be essentially impossible to recognise. This is especially true for message receivers who may not possess visual recognition capabilities, at least not visual capabilities that are similar to our own vision.
Refer to caption           Refer to caption
Figure 4: Left: The original Arecibo message intended to be reconstructed, but sent as a linear stream from the radiotelescope in Arecibo, Puerto Rico. The 1,679 bits are meant to be arranged into 23 columns of 73 rows, 23 and 73 being two prime numbers which when multiplied together equal 1,679. Right: If the stream is instead arranged into 23 rows and 73 columns, the original visual interpretations of the message are scrambled, which may result in a figure that is closer to being statistically random. What we show is that the message is still there, concealed, and can be deciphered by algorithmic deconvolution.
Figure 5: Top left: Most possible partitions result in random-appearing configurations with high corresponding complexity, indicating measurable randomness. Bottom: Some partitions will approximate the originally encoded meaning (third from the right). Other configurations result in images with higher complexity values. This sequence of images shows the images in the approximate vicinity of the correct bidimensional configuration (i.e., partition) and illustrates fast convergence to low complexity. Top right: By using different information indexes across different configurations, a downward-pointing spike will indicate message (image) configurations that correspond to low-complexity image(s). This allows a prior-knowledge-agnostic and objective method to infer a message’s original encoding. Of the various measures, BDM, combining classical information (entropy) for long ranges and a measure motivated by algorithmic probability for short ranges, is the most sensitive and accurate in this regard. Traditional compression and entropy also indicate the right configuration amongst the top spiking candidates. The ratio of noise-to-signal was amplified in favour of the hidden structure by multiplying the original image size by 6 for both length and height.
In addition, this message was reformatted from its original bidimensional state and sent as a string of 1,679 binary digits. The reasoning was that an alien civilisation with a mathematical system that receives the message will recognise 1,679 as a semi-prime number–a multiple of 23 and 73. Even if the original message could not be reconstructed in its original and intended encoding, the receiver would recognise the message as having some mathematical significance that is unlikely to be a random signal from outer space.
Fig. 5 shows how the method provides a robust approach to identifying the best way to decode a signal back to its original multidimensional space, by sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages by taking the lowest algorithmic complexity configurations under the assumption that the original message is algorithmically not random (i.e., that the original message is compressible). As shown in [29] for graphs, this is because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very low. On the other hand, if the original message is sufficiently compressible, and therefore not random, most partitions will lead to configurations of greater algorithmic randomness.
Figure 6: The method’s resilience in the face of some noise. At 3% of the bits of the original $23\times 73=1679$-pixel image randomly flipped (which means about 1.5% were binary negated), the method remains sensitive and displays a small downward spike at the 23 value, uncovering its length, but the signal gets lost when more bits are flipped.
Fig. 6, shows the method’s noise resistance to indicate the candidate partition and the precise lengths of the original message. Successful amplification of the signal is shown in Fig. 7, with BDM outperforming Compression and Shannon Entropy in the face of additive noise, with Compression showing insensitivity to the original signal at about 10% of bits flipped, and Shannon Entropy diminishing faster than BDM but slower than Compression. In all cases, different levels of robustness at deconvolving the image dimensions are shown.
Investigating this further, Fig. 9 shows six bidimensional images along with their original numerical dimension. The values of the three complexity measures over possible partitions are shown below the original numerical dimension for ease of comparison. The method is invariant with regard to linear transformations such as encoding, as shown in images 2 and 4 of Fig. 9.
Drops in complexity (downward spikes) indicate candidate dimensions for the original encoding dimensions, and thus the decoding dimension that would process the signal such that it produces the lowest-complexity message. BDM outperforms compression and entropy and correctly identifies the original message (image), encoding for 4 out of 6 images.
Figure 7: Applying 3 different quantitative methods, it is shown (left) that they are highly insensitive to signal and highly sensitive to noise (16.5% of pixels randomly flipped). However, by growing the original image (right) by a factor of 6 on each dimension (i.e. 1 pixel becomes $6\times 6$), the methods are less sensitive to noise and more sensitive to signal, with BDM significantly outperforming Compression, and Shannon Entropy showing sensitivity at up to 60% pixels flipped (hence about 30% of the original image) versus Compression and Shannon Entropy, that are about 50% sensitive. Downward spikes (right) are shown at $23\times 6=141$.
Figure 8: Sequence of lowest complexity (BDM value on top in bits) partitions of a black and white image of a space shuttle, with the top right partition as the correct reconstruction and other candidates cases in which an alignment occurs at some multiple of the correct dimension.
Some of these image decodings for the space shuttle image are shown in Fig. 8. Each panel shows random realignments of the correct dimensions along with the original image (i.e., object) embedded into these new partitions, with their respective BDM values displayed above each image. When compared to the other bidimensional spaces, the one that is most similar to the original message will have a lower complexity value.
By translating (encoding) a colour image into binary black and white pixels, this method picks an image decoding that is similar enough to the original image (at least similar enough visually). Fig. LABEL:mandril shows the original image on the left, the results of the method on the right, and the resulting image selected from the spike in BDM on the bottom. A mirror-like image is selected as the first candidate, followed by the correct one. This kind of spiking also suggests the original partition can be inferred via smaller, non-random spikes at multiples or divisors of the native partition (500). In this case, small spikes at half the original partition (250) provide clues to the original embedding in 2D.
196                                   82                                     215
276                                 360                                 220
Figure 9: Six 2D images (labelled 1 through 6 going from left to right, top to bottom) of very different nature, including a demonstration of linear-transformation invariance conforming to the underlying theory (in this case, rotation of a mathematical formula). Size invariance has actually shown amplification of signal-to-noise difference. Under each image is its correct numerical (first) dimension. The values of the three complexity indexes over possible partitions are shown below the original numerical dimension for ease of comparison. Downward spikes indicate candidates for possible original partitions. In all cases, the correct dimension value is among the top three candidates, with BDM outperforming in 4 out of 6 cases at indicating the top candidate.
A.5 Further discussion
A.5.1 Algorithmic information dynamics in communication problems
Algorithmic information dynamics (AID) [24, 25, 26] consists of a combination of perturbation analysis and both classical information theory and algorithmic information theory in order to solve inverse problems and perform causal analysis by investigating information and randomness dynamics. It searches for possible generative models, which result from perturbations, and tests whether or not they are compatible with the gathered (or intended) phenomena to be explained (or predicted), while this act of sweeping does not in turn influence the generative models found (i.e., it is a blind search). The underlying idea is that a computable model is a causal explanation of a piece of data. These models can later be exploited as testable underlying mechanisms and causal first principles [25, 28], regardless of those being stochastic, computable, or mixed processes.
Such an investigation process is based on the universal optimality of algorithmic probability proved in AIT so that algorithmic complexity values are approximated and combined with Shannon entropy by using the block decomposition method (BDM). Thus, AID constitutes a formal approach to Artificial General Intelligence that requires a massive production of a universal distribution, the mother of all models [27], to build a very large (semi-computable) model of computable approximate models. The semi-computability of algorithmic complexity (or, equivalently, of the universal distribution) means that one can approximate its exact value from above but never with a known and specified precision.
Our theoretical results in Section A.3 show that the method employed to achieve the empirical findings in Section A.4 is sound with regard to the infinite asymptotic limit when computational resources are unbounded; and in addition it also shows the method can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Thus, our empirical results suggest that AID is capable of harnessing (under feasible computational resources) some of the algorithmic-informational properties that we demonstrated in ideal and generalised scenarios. As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, and most prominently for high algorithmically probable objects.
As explained in Section A.1, the (algorithmic) perturbations that we investigate in this work can change either the object (e.g., by flipping bits) or its multidimensional space (e.g., by reconfiguring the dimensions’ lengths of a given partition to achieve another partition) independently, or can change both at the same time. The reader is invited to notice that the present work applies AID to the problem of communicating messages by performing a perturbation analysis on the received signal. Thereafter, the principles of AIT are applied through approximations to algorithmic complexity by using BDM. These approximations are employed as a performance measure that characterizes the criterion to estimate the correct multidimensional space by distinguishing it from the wrong ones, which are the outcomes of the perturbations. Therefore, as also raised in Section A.2, our present approach, or AID in general, cannot be reducible to a straightforward application of AIT (e.g. by using the algorithmic coding theorem to approximate the optimal distribution or another method to approximate algorithmic complexity [19]) so as to estimate the irreducible information content of the received signal in order to directly retrieve the original message sent by the emitter agent.
In fact, it is the semi-computability that evinces the divergence power brought about by the landscape of the approximate complexity values, landscape which results from the perturbations and from which we can find the best candidates to choose. One could empirically estimate the values of the algorithmic complexities of each possible message and also compare them with each other in order to check which one displays the lowest algorithmic complexity, therefore more likely to be the correct one. However, due to the uncomputability of algorithmic complexity (or algorithmic probability), this approach alone will never assure that the estimation error is not necessarily high enough so as to make the complexity value of a message actually be lower than the complexity value found in another message, whose approximate complexity value initially displayed a lower value in comparison to that of the first message in the empirical estimation process. In other words, the uncomputable errors in any estimation of algorithmic complexity can deceive one into thinking a message which actually has higher complexity than another message, has a lower complexity than the latter. The objects in other dimensions may not actually be more random, they only appear more random to the approximating methods, but it is the perturbation analysis that helps us to find the intended patterns.
Although it may seem paradoxical at first glance—because one may initially think that the errors produced by the approximating methods to the algorithmic complexity would be inherited by the perturbation analysis phase itself—, it is the perturbation analysis that enables one to overcome such a semi-computability limitation but only when looking at the whole landscape of complexity values one estimates with a sufficiently large amount of perturbations effected on the received signal. Enabled by the employment of AID in the problem of interpreting a signal without prior knowledge of its source, we believe this also is one of the contributions of our results. As the number of perturbed messages strongly dominate the number of possible best message candidates, our theoretical and empirical results show that perturbation analysis enables the downward spikes displayed in the landscape of complexities to more likely indicate the best message candidates, even though the exact values of the algorithmic complexities are uncomputable and one can only asymptotically approximate them from above.
The underlying reason this occurs is because the algorithmic complexities of the messages resulting from the algorithmic perturbations eventually become sufficiently larger than the algorithmic complexity of the original message. That is, the set of more random candidates becomes much larger than the set of compressible candidates. Hence, a powerful enough approximating method to the algorithmic complexity, such as BDM, can harness this divergence in order to enable the perturbation analysis to statistically set apart the correct one (i.e., more compressible) from the wrong ones (i.e., more random). This holds because an approximating method with enough computational resources can produce smaller errors than the actual difference of complexity values between the correct candidate and the perturbed ones, even though the exact complexity values remain ultimately unknown to the perturbation analyst (i.e., even though the analyst itself can never be sure it has enough computational resources in the first place). Individually, the approximation to the complexity value of each object and multidimensional space is prone to the unknown error margins that follow from the semi-computability. However, because each pair object-space results from a perturbation which is a particular case of algorithm and each algorithm theoretically has an exact complexity value associated to it, one can retrieve a likely correct answer from a sufficiently large landscape of these perturbations.
Thus, although there is an intrinsically subjective “deficiency” in the perturbation analysis that forbids an arbitrarily precise approximation to algorithmic complexity, this subjective aspect can be eventually cancelled out in the perturbation analysis phase. In case one chooses another complexity measure that is in principle computable instead of the measures given by algorithmic information theory, to be employed in the analysis phase, then Equation 25 would not hold in general, and in principle [18, 19]. Therefore, the subjective aspects, such as the arbitrarily chosen programming language, probability measure $\mathbf{P}$, or the computational appliances made available, would not be able to be overcome in the perturbation analysis phase. In this sense, our results show that it is the very subjective “deficiency” given by the semi-computability that counter-intuitively enables the perturbation analysis to eventually overcome this and the other subjective aspects, that is to find an objective piece of information.
A.5.2 Toward an algorithmic noisy-channel coding theorem
The reader is invited to notice that Theorem 11 demonstrates an algorithmic-informational version of the noisy-channel coding theorem in classical information theory [20]. Instead of the purely stochastic noise to which the communication channel may be subject in the classical version, the communication problem we studied in this paper deals with communication channels under algorithmic noise, i.e., the case in which signals are prone (or may be subject) to algorithmic perturbations [26] that can change the originally sent signals like a computer program could do, then resulting in new signals—which may substantially differ from the respective original signals. This new version of the problem of communicating a message can be understood as being more general than the classical version because: it holds for zero-knowledge one-way555 Notice that the classical version also applies to one-way communication but since both agents must know each other’s encoding-decoding scheme beforehand, one may argue that at least some information was previously communicated between the emitter agent and the receiver agent, which by definition violates the condition of one-way communication in the first place. communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26].
However, in addition to Theorem 11 requiring unbounded computation time so that one can calculate the optimal values of algorithmic complexity, one in fact only proves in Theorem 11 sufficient conditions for enabling the (one-way) communication in channels under algorithmic perturbations (i.e., under algorithmic noise, which subsumes but differs from the stochastic noise in the classical version) and zero knowledge about the encoding scheme chosen by the emitter. Instead, for a complete analogous to the classical version, one still needs to demonstrate necessary conditions, that is, to investigate tighter bounds for the asymptotic dominances regarding the relationship between the algorithmic complexity of the algorithmic perturbations and the algorithmic complexity of the message itself. Thus, while we have demonstrated sufficient conditions so that an algorithmic noisy-channel coding theorem can hold, future research is paramount for investigating its necessary conditions, like achieving an algorithmic analogous to the channel capacity, which in the classical version plays the role of a threshold for enabling or not the communication (with errors as low as one wishes).
A.5.3 Zero-knowledge communication and ZKPs
As shown in the main article and discussed in the previous Section A.5.2, one of the salient advantages of our results in Theorem 11 is the absence of a condition stating that the encoding-decoding scheme arbitrarily chosen by the emitter agent must be also known by the receiver agent. Such a condition that relaxes the need of prior knowledge is what defines the zero-knowledge characteristic of the problems we investigated in this work. Our results are not only agnostic to prior knowledge of encoding-decoding schemes used to encode the message by the emitter agent, but also to: the computation model and the programming language chosen in order to implement our proposed method; the computable (or semi-computable) method of approximation to algorithmic complexity; any arbitrarily chosen (computable) probability measure of the events; and the formal theory able to formalize the mathematical statements. We have demonstrated sufficient conditions in this zero-knowledge scenario for enabling the reconstruction of the original message (i.e., the original object embedded into the original multidimensional space to be conveyed to the receiver) in one-way communication channels.
In this manner, zero-knowledge communication (ZKC) occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of no prior information about the emitter agent only applies to encoding-decoding schemes (and programming languages and computation models), and consequentially also to the original multidimensional space and object that are yet unknown by the receiver agent before any communication takes place. However, this zero-knowledge condition does not mean other assumptions regarding the unknown emitter are not being considered by the receiver (in addition to those mathematical conditions already investigated in Section A.3). For example, as it is the case studied in the present work, the assumption that the emitter agent is somehow capable of performing an arbitrary encoding and compression of the message into a signal stream.
Not only because of the nomenclature, but also because of a mathematical proposition being ascertained with probability arbitrarily close to $1$ as a larger number of events happen, one may be tempted to relate ZKC with zero-knowledge proof (ZKP) as traditionally known in cryptography [21, 22, 23]. Indeed, one may consider that they share some features or intuition in common. For example, notice that the one-way characteristic in ZKC problems plays the role of a direct analogous to non-interactive ZKPs. Nevertheless, it is important to also notice that ZKP and ZKC are based upon opposing assumptions and goals regarding the “zero-knowledge” characteristic.
In the ZKP case, one looks for high enough probabilities of one agent correctly attesting the other agent’s achievement while no information (or knowledge) is revealed about the actual content of this achievement (e.g., the proof of a theorem). Intuitively, one agent (i.e., the prover) should statistically convince the other agent (i.e., the verifier) of having certain information (i.e., knowledge) in the long run, and this convincing should happen in such a way that the second agent acquire no information about what the former managed to convince the latter.
In the ZKC case instead, one looks for high enough probabilities of one agent being correct about the very content of what the other agent intended to communicate, given that no information was known by the former about the latter. Intuitively, one agent (i.e., the receiver agent) should statistically guarantee that its own interpretation of the signal sent by the other agent (i.e., the emitter agent) to the former corresponds to the original message that the latter intended, and this understanding (i.e., decoding) should happen in such a way that no prior information (i.e., knowledge) about what and how the second agent will communicate in the first place is known.
On the one hand, the resulting process in ZKPs should lead to no knowledge being acquired as a goal, once the acquisition of any knowledge is undesirable. On the other hand, the resulting process in ZKC should lead to maximum knowledge being acquired as a goal, once zero knowledge was a starting condition in the first place. In this sense, one may consider ZKP and ZKC as akin mathematical problems, but diametrically opposing counterparts of each other regarding the acquisition of knowledge.
A.5.4 A perturbation analysis-based method for the maximum algorithmic likelihood
In order to tackle the new challenges presented in Section A.2, we have investigated additional conditions that enable one to infer the correct message, given that the receiver has no prior knowledge about the source (i.e., object and multidimensional space). In this sense, one aspect of the contributions of the present work resides in demonstrating the existence of sufficient conditions that allows the application of the algorithmic coding theorem [39, 41] to estimate the best model or parameters, even in scenarios where algorithmic information distortions may take place due to isomorphic transformations of the multidimensional space. Being one of the central results and powerful tools from algorithmic information theory, the usage of the universal distribution established by the algorithmic coding theorem underpins methods based on universal (Solomonoff) induction or minimum description length  [27, 40], whose general solutions encompass and refine the maximum likelihood or maximum entropy methods [20, 40, 29].
As discussed in Section A.5.1, another aspect of our empirical and theoretical results resides in showing how perturbation analysis can be employed to overcome some of the semi-computability limitations in the approximations to algoritmic complexity (or algorithmic probability) values. Because the algorithmic complexities of the perturbations dominate those of the best candidates, one can still statistically infer the most likely messages (associated with higher algorithmic probabilities) due to the increasing divergence between the actual complexity values of these best candidates and the perturbed signals. By taking the algorithmic complexity of the object and the multidimensional space as the parameters to be estimated, our methods and results demonstrate how one can statistically infer the parameters with highest algorithmic likelihood [33, 34] from perturbation analysis without the need of exactly calculating the (uncomputable) algorithmic complexity values. In this regard, as demonstrated in Theorem 11, this inference method is general enough to remain robust vis-a-vis any arbitrarily chosen alternative method to calculate the probabilities of the events, while assuming no bias in the priors (i.e, while taking a uniform distribution of the possible messages as starting point).
For example, Theorem 11 shows that for any probability distribution, encoding-decoding scheme, programming language, and computation model, the algorithmic complexity of the correct message is strongly dominated by the algorithmic complexities of the perturbations, dominance which holds with (uniform) probability arbitrarily close to $1$. If one knows how to calculate an upper-bound approximation to the algorithmic complexities (approximations which can be calculated by computable methods), then there are sufficient conditions for which one can in general infer that the algorithmic complexity of the sent message increases much slower than the algorithmic complexity of the perturbations, or equivalently that the algorithmic probability of the sent message increases much faster than algorithmic complexity of the perturbed signals.
Intuitively, as it is already known to hold analogously for Fisher information [20] (which gives a lower bound for the variance of any unbiased estimator) in traditional statistics, sharper and more isolated valleys (i.e., downward spikes) found in the landscape of complexity values calculated for each parameter should indicate a better approximation to the actual algorithmic complexity (or algorithmic probability) of the correct parameter. Formally, if the algorithmic complexities of the perturbed signals are known to be sufficiently higher than the approximations to the complexities of the best message candidates (along with the fact that the quantity of distinct perturbations also strongly dominates the quantity of best candidates), then Theorem 11 shows that the errors in these approximations are strongly dominated by (i.e., increase much slower than) the highest algorithmic complexity of the perturbations. Thus, finding the existence of more tighter bounds in the relationship between the variance of the algorithmic complexity change rate (e.g., a rate obtained by calculating the ratio between the information difference [37] and the algorithmic complexity of the perturbation that changed the elements of the system) and how small those errors are is a problem that demands future investigation. We believe that our present work also contributes by establishing first results that posit a new path of future research in this direction.